### v1

# On the resolution of categorical symmetries in (Non-) Unitary Rational CFTs

Abstract
We explore several aspects of the categorical symmetry-resolved entanglement entropy (SREE) in two-dimensional Rational Conformal Field Theories (RCFTs) and express it directly in terms of the modular data of the theory. Motivated by the SymTFT construction of Choi:2024wfm, we provide a general formula for SREE that applies to symmetric (weakly/strongly) and cloaking boundary conditions as well as for fusion rings with multiplicities. We check the formula against several explicit examples. Additionally, we study symmetry resolution for both categorical and invertible symmetries in (non-)diagonal RCFTs and comment on the subtleties that arise in these cases. Finally, we extend our analysis to diagonal non-unitary RCFTs, focusing on theories with generalized Haagerup–Izumi modular data, and find full agreement with the given formula.
1 Introduction
Symmetries are omnipresent in nature. In physics, symmetry plays a crucial role in understanding the dynamics of quantum field theory and provides a framework for going beyond. According to Wigner Wigner:1959, each symmetry is associated with a unitary operation that has an inverse, i.e., group-like. Gaiotto et al. Gaiotto:2014kfa showed that the generalized global symmetry of a theory is associated with a topological defect operator, and there are plenty of examples found in condensed matter systems that realize this generalized global symmetry McGreevy:2022oyu. One of the classes of these generalized symmetries is famously known as non-invertible symmetry Bhardwaj:2017xup; Chang:2018iay; Thorngren:2019iar; Heidenreich:2021xpr; Thorngren:2021yso; Choi:2022jqy; GarciaEtxebarria:2022vzq; Cordova:2022ieu; Cordova:2022ruw; Choi:2022fgx; Choi:2022rfe; Apruzzi:2022rei; Shao:2023gho; Seiberg:2023cdc; Schafer-Nameki:2023jdn; Bhardwaj:2023wzd; Bhardwaj:2023ayw; Bhardwaj:2023fca; Choi:2023pdp; Lin:2023uvm; Benedetti:2023owa; Iqbal:2024pee; Choi:2024rjm; Seiberg:2024gek; Bhardwaj:2024wlr; Bhardwaj:2024kvy; Putrov:2024uor; Shao:2025mfj; Shao:2025qvf  [^1] The references mentioned here is by no means exhaustive. Interested readers are referred to some of these reviews and lecture notes and their citations McGreevy:2022oyu; Shao:2023gho; Schafer-Nameki:2023jdn for more references., which can not be described by a unitary operation, more specifically, it does not have an inverse. In mathematics literature, it is known as a category Bartsch:2023wvv; Costa:2024wks. For these kinds of categorical (non-invertible) symmetries, associated topological operators $\{\mathcal{L}_{a}\}$ obey fusion rules of a tensor category rather than a group law, as shown below,
$\mathcal{L}_{a}\times\mathcal{L}_{b}={\bigoplus}_{c}\mathcal{N}_{ab}^{\,\,\,c}\,\mathcal{L}_{c}$ (1)
where, $\mathcal{N}_{ab}^{\,\,c}\in\mathbb{Z}_{\geq 0}$ is the fusion coefficients. These defect lines usually do not have a formal inverse, i.e., there exists no $\mathcal{L}^{-1}$ such that $\mathcal{L}\times\mathcal{L}^{-1}=\mathbb{I}$.
There are numerous examples that demonstrate this kind of categorical symmetry exists in various branches of physics, including high-energy theory, mathematical physics, and condensed matter physics, particularly in low-dimensional quantum systems that are useful for quantum computation Preskill:1999he; Kitaev:2005dm; Kitaev:1997wr; PhysRevLett.96.110405; Aharonov:2005nkd; Feiguin:2006ydp; Dennis:2001nw; McGreevy:2022oyu; PRXQuantum.5.037001; Okada:2024qmk; Li:2024gwx; Bhardwaj:2024ydc; PhysRevX.15.011058; Li:2025bgo; KNBalasubramanian:2025vum. Before delving into the main discussion of the paper regarding the categorical symmetry resolution of the entanglement entropy, we briefly discuss the representation of categories and the importance of dealing with the Drinfeld centre or the quantum double $Z(\mathcal{C})$ of a given category $\mathcal{C}\,.$
Fusion category and tube algebra.
As discussed above, the symmetry operations will be encoded by a fusion (tensor) category $\mathcal{C}$ generated by topological lines $\{\mathcal{L}_{a}\}$. While $\mathcal{C}$ is not itself an algebra, one can canonically associate a tube algebra to it using the construction of Ocneanu Ocneanu2001 as follows. Consider an annulus (“tube”) whose inner and outer boundaries are labeled by $i,j\in\mathrm{Irr}(\mathcal{C})$ and with a line $x\in\mathrm{Irr}(\mathcal{C})$ running along the tube’s axis. Stacking such tubes and recoupling via $F$-moves defines a product on the space
$\mathrm{Tube}(\mathcal{C})=\bigoplus_{x,i,j\in\mathrm{Irr}(\mathcal{C})}\mathrm{Hom}_{\mathcal{C}}(i\!\otimes\!x,\,x\!\otimes\!j),$ (2)
making it a finite-dimensional *-algebra; the involution is given by orientation reversal (Hermitian adjoint on morphisms). There is a canonical equivalence cite-key; 2001math.....11205M
$\texttt{Rep}\,\big(\mathrm{Tube}(\mathcal{C})\big)\;\simeq\;Z(\mathcal{C})\,.$ (3)
Therefore, simple $\mathrm{Tube}(\mathcal{C})$–modules are in one-to-one correspondence with the simple objects (the “anyon types”) of the Drinfeld center $Z(\mathcal{C})$. In other words, the irreducible representations of $\mathrm{Tube}(\mathcal{C})$ are canonically identified with the simple topological defect lines of $Z(\mathcal{C})$, and each representation is characterized by the quantum dimension of the corresponding simple object. Consequently, for a quantum system whose categorical symmetry is described by a semisimple fusion category $\mathcal{C}$, the analysis of its symmetry sectors is naturally equivalent to studying the symmetry sectors of the Drinfeld center $Z(\mathcal{C})$.
Categorical symmetry–resolved entanglement.
Parallel to recent advances in the characterization of (non-)invertible symmetries, a complementary line of research has focused on the role of symmetries in shaping entanglement structures in quantum field theories (QFTs). Entanglement serves as a fundamental measure of quantum correlations and phases of matter, and provides a particularly sensitive probe of underlying symmetry principles, including generalized and non-invertible symmetries. In this context, symmetry-resolved entanglement entropy (SREE) Goldstein:2017bua; german3; Belin:2013uta; Islam:2015mom; doi:10.1126/science.aau0818; FG; Murciano:2019wdl; Murciano:2020lqq; Parez:2021pgq; Murciano:2020vgh; Zhao:2020qmn; ourPartI; Horvath:2021rjd; Ares:2022gjb; DiGiulio:2022jjd; Gaur:2023yru; Kusuki:2023bsp; Northe:2023khz; Gaur:2024vdh; Banerjee:2024ldl [^2] For more references and details, interested readers are referred to the following review of symmetry-resolved entanglement measure Castro-Alvaredo:2024azg. has emerged as a refined diagnostic that decomposes the total entanglement into contributions associated with individual symmetry sectors, thereby elucidating how symmetry structures govern the distribution of quantum information. Recent developments further demonstrate that non-invertible symmetries can impose nontrivial constraints on entanglement patterns, influence topological order, and enforce selection rules that transcend those encountered in the setting of ordinary invertible symmetries Casini:2019kex; Copetti:2024dcz; AliAhmad:2025bnd. Motivated by these developments, this paper examines how non-invertible (categorical) symmetries modify entanglement measures in certain two-dimensional topological systems at criticality. For global invertible symmetries, SREE has been extensively studied, allowing the entanglement to be resolved sector by sector. Readers are referred to these reviews Castro-Alvaredo:2024azg. In contrast, symmetry resolution for categorical symmetries Lin:2022dhv; Sierra_2024; Das:2024qdx; Heymann:2024vvf; Choi:2024wfm; Das:2025xyz remains far less explored. In (1+1)(1{+}1)-dimensional CFTs with a categorical symmetry encoded by a fusion category $\mathcal{C}$, symmetry resolution can be formulated in terms of boundary CFT (BCFT) DiGiulio:2022jjd; Kusuki:2023bsp; Northe:2023khz through projections onto categorical sectors. These projections are implemented by symmetric (either strongly or weakly) boundary states on which the corresponding topological defect lines can end. Now, on the replica geometry Calabrese:2009qy, inserting compatible defect endpoints along the branch cut yields well-defined sector probabilities and sector-resolved Rényi and von Neumann entropies. This construction is valid whenever the categorical symmetry is non-anomalous, i.e. free of obstructions to gauging Choi:2023PRD; Seifnashri:2025fgd. Alternatively, there is a twist-field approach that compute directly symmetry-resolved entanglement entropy using charged moments and this method is applied to compute SREE for rational CFT with Lie group symmetry by Calabrese:2021wvi and for certain non-invertible Das:2025xyz.
Interestingly, for 2d CFTs with global internal abelian symmetry, SREE shows universal results in the sense of equipartition of entanglement among all charge sector german3; Castro-Alvaredo:2024azg. On the other hand, in the context of rational CFTs (RCFTs) [^3] It turns out that almost all rational CFTs possess non-invertible symmetry., for non-invertible categorical symmetry-resolved entanglement entropy, this equipartition of the entanglement is broken Kusuki:2023bsp. These deviations are controlled by the fusion data of $\mathcal{C}$ and by the choice of symmetric boundary condition. Note that computation of the entanglement requires partitioning the Hilbert space, hence the boundary state placed at the entangling surface plays a crucial role in refining the entropy Ohmori:2014eia; Cardy_2016. In contrast to the group-like setting, non-invertible symmetries generally admit multiple, and often inequivalent, choices of boundary conditions. Although this issue has been noted in the literature Choi:2024wfm; Choi:2023PRD; Choi:2024tri; Lin:2022dhv; Heymann:2024vvf; Das:2024qdx, previous discussions have largely centered on the so-called weakly/strongly symmetric boundary condition and with a multiplicity-free fusion ring. In this work, we revisit this question through a series of concrete examples that clarify how different boundary conditions naturally arise and how they affect the symmetry resolution. In particular, we highlight situations in which the weakly symmetric boundary state is absent and motivate the use of cloaking boundary conditions (or states) Brehm:2022JPA; Brehm:2024arXiv, which, despite not corresponding to physical Cardy states, play a crucial role in implementing symmetry resolution in non-invertible symmetries consistently.
Non-unitarity and symmetry resolution.
Most developments on symmetry resolution have focused on unitary settings, where the physical Hilbert space and defect operations admit a positive-definite inner product, with positive central charge. In contrast, symmetry resolution for categorical symmetries in non-unitary setting remains comparatively unexplored, despite clear motivation. By the bulk–boundary correspondence, when a (2+1)(2{+}1)-dimensional topological phase is described by a unitary TQFT, the universal properties of its gapless boundary are captured by a two-dimensional unitary RCFT Witten:1988hf; Moore:1988qv; Moore:1989vd; Dijkgraaf:1989pz; Verlinde:1988sn; Fuchs:2009iz. Extending this logic beyond unitarity is natural Gang:2023rei. It has been argued that a family of non-unitary TQFTs arises from topological twists of certain three-dimensional $\mathcal{N}=4$ rank-zero (S-fold) SCFTs, producing modular data that generalize the non-unitary Haagerup–Izumi modular data Gang:2023ggt; Gang:2022kpe. This, in turn, leads to the question of whether there exist boundary RCFTs compatible with such generalized Haagerup modular data Haagerup1994; Evans:2010yr. The answer is affirmative: using the so-called Riemann–Hilbert approach, conformal characters consistent with the proposed modular data have been explicitly constructed Gang:2023ggt. These developments supply a concrete arena in which to formulate and study categorical symmetry resolved entanglement for non-unitary RCFTs, thereby extending symmetry resolution beyond the unitary paradigm. At this point, note that there are recent studies aimed at constructing critical lattice models in two dimensions, realizing the symmetry associated with the unitary Haagerup fusion category $H_{3}$ Huang:2021nvb; Vanhove:2021zop; Jia:2024wnu; Bottini:2025hri and extracted the central charge (and operator spectra) which turns out to be $c\sim 2$. Then it was further conjectured that this two-dimensional CFT can be constructed by a taking coset of a Haagerup rational CFT characterized central charge multiple of $8$ Evans:2010yr, even though there is a puzzle regarding the existence of such $c=8$ unitary rational CFT with Haagerup fusion $\mathcal{Z}(H_{3})$ (for more details, interested readers are refered to T. Gannon’s talk Gannon:2025birs) [^4] Recently, entanglement asymmetry for $\mathcal{Z}(H_{3})$ has been computed Benini:2025lav.. Nonetheless, we draw motivations from these studies and focus on certain non-unitary generalized Haagerup fusion categories Gang:2023ggt(some of them are tensor products of certain non-unitary minimal models) and discuss symmetry-resolution of entanglement entropy.
Goal and brief takeaway.
Besides the mentioned developments in the computation of SREE, the resolution in non-diagonal and non-unitary theories remains immensely unexplored. In this paper, we intend to fill this gap. This paper aims to bridge this gap by extending the analysis of categorical symmetry resolution to the domains of non-diagonal and non-unitary RCFTs. We provide a general computational framework for computing the SREE in these theories, expressing it directly in terms of their modular data. Our work provides a unified framework that encompasses a wide range of scenarios, including different types of boundary conditions and fusion rings with multiplicities. We test the general formula against several explicit examples, including unitary diagonal models such as Ising CFT, tetracritical Ising CFT, and non-unitary models realized by generalized Haagerup-Izumi modular data. We also make comments on the subtleties and limitations of BCFT techniques while applied to non-diagonal theories with the example of non-diagonal $SU(2)_{10}$ type $E_{6}$, and its $\tfrac{1}{2}E_{6}$ resolution.
Our paper is organised as follows. In Section (2) we review the essential ingredients and terminology used throughout the work. We outline the different types of boundary conditions relevant for our analysis. Section (3) presents the general strategy for computing symmetry-resolved entanglement entropy (SREE) by decomposing the entanglement into contributions from individual representations. We examine how different choices of boundary states inserted at the entangling surface of a bipartite system influence the resolution of entanglement. In this context, motivated by the SymTFT construction of Choi:2024wfm, we provide a general formula for SREE that applies to a broad class of theories, including non-unitary and non-diagonal RCFTs, and accommodates various admissible boundary conditions. In Section (4) we compute SREE in several unitary models and verify the consistency of the corresponding boundary entropies using the generalized Affleck–Ludwig boundary entropy formula. We also attempt to resolve the symmetry sectors of the non-diagonal $SU(2)_{10}$ WZW model with $E_{6}$ type classification, and discuss the limitations of BCFT techniques in resolving its $\frac{1}{2}E_{6}$ categorical symmetry. Nevertheless, we demonstrate that such limitations are not universal: for example, we explicitly resolve the $\mathbb{Z}_{3}$ symmetry (an invertible symmetry) in the three-state Potts model, which stands as a non-diagonal RCFT that admits consistent symmetry resolution using BCFT techniques. Finally, in Section (5) we turn to several non-unitary models and show that the modified Affleck–Ludwig formula continues to hold in these cases as well, thereby extending our analysis beyond the domain of unitary RCFTs. We conclude with a discussion in Section (6).
2 Background and basic setup
We begin by reviewing the fundamental principles and structural properties of RCFTs. An RCFT is distinguished by the existence of only finitely many primary fields together with a rational central charge $c$. A central role in their formulation is played by the construction of the Vertex Operator Algebra (VOA), which provides the algebraic foundation for the theory. In general, the partition function of an RCFT takes the form DiFrancesco:1997nk,
${Z}(\tau,\,\bar{\tau})=\sum_{i,j}M_{ij}\,\chi_{i}(\tau)\bar{\chi}_{j}(\bar{\tau})\,\,\,\,\,\,\,\,\,\,\,\,\,i,j\in\textrm{Primary spectrum}$ (4)
where, the characters $\chi_{i}(\tau)$ are defined as,
$\chi_{i}(\tau)=\text{Tr}_{\mathcal{H}_{R_{i}}}\,q^{L_{0}-\frac{c}{24}}\bar{q}^{\bar{L}_{0}-\frac{c}{24}},\textrm{with,}\,\,\,q=e^{2\pi i\tau}$ (5)
along with the Hamiltonian,
$H=L_{0}+\bar{L}_{0}\,.$ (6)
Now the degeneracy matrix $M_{ij}$ in (4) need not be diagonal always. When $M_{ij}=\delta_{ij}$, we call the RCFT a diagonal one; otherwise, we call it a non-diagonal RCFT. In the next subsection, we will briefly introduce the notion of boundaries and defects, which will be relevant for our computation of SREE in the subsequent sections.
Introducing defects and conformal boundary conditions in RCFT
Figure 1: Inserting a topological defect line($\mathcal{L}$) along the Euclidean time direction (left), which, after the conformal transformation (from cylinder to plane), acts on the local operator (right).
A conformal defect can be understood as an extended operator that extends along the Euclidean time direction while remaining localized in the spatial dimensions Shao:2023gho; Schafer-Nameki:2023jdn as shown in Fig. (1). For non-invertible defects, the corresponding defect lines are known to be equivalent to the charge operators in the associated invertible theory. This correspondence offers a conceptual framework for computing the SREE through the evaluation of charge moments Goldstein:2017bua; german3; Bonsignori:2020laa. In contrast, our work adopts a different strategy: we employ the established BCFT formalism to determine the SREE Kusuki:2023bsp. Regardless of the chosen approach, the SREE can ultimately be expressed in terms of the quantum dimensions of the defect lines contributing to each symmetry sector. For the construction of the boundary states, we choose the conformal Cardy states. These cardy states are the linear combination of the Ishibashi states with specific coefficients. Ishibashi states are solutions of the Ishibashi condition given by the construction of BCFT as Ishibashi:1988kg; Cardy:1986gw; Cardy:1989ir; Cardy:1989vyr; Cardy:1991tv; Cardy:2004hm  [^5] For more details, we suggest the reader the following review Northe:2024tnm and the book Recknagel:2013uja. ,
$(T(z)-\bar{T}(\bar{z}))|_{\rm{bdy}}=0.$ (7)
Now using Virasoro mode decomposition, $T(z)=\sum_{n=0}^{\infty}\frac{L_{n}}{z^{n+2}}$ for the Ishibashi condition, we get the following equation,
$(L_{n}-L_{-n})|\alpha\rangle\rangle=0$ (8)
where, $\{L_{n}\}$ are the Virasoro generators.
Strong and Weakly symmetric boundary: We define a boundary weakly symmetric if the symmetry defects can end topologically on it, leading to conserved operators for the Hamiltonian on a strip (open string channel). Similarly, for a boundary becoming strongly symmetric, the relevant boundary state is an eigenstate of the symmetry operators (closed string channel). These two notions are degenerate for invertible symmetries but quite different for non-invertible symmetries. We define them as follows. The boundary is called strongly symmetric if for every defect line $\mathcal{L}_{a}$ for $a\in\mathcal{C}$ there a boundary state $|\!|B\rangle\!\rangle$ which satisfies the following Choi:2023PRD; Choi:2024tri,
$\mathcal{L}_{a}|\!|{B}\rangle\!\rangle=\langle\mathcal{L}_{a}\rangle|\!|{B}\rangle\!\rangle\hskip 85.35826pt\forall a\in\mathcal{C}$ (9)
and similarly a boundary is called weakly symmetric if the following equation holds,
$\mathcal{L}_{a}|\!|{B}\rangle\!\rangle=|\!|{B}\rangle\!\rangle+\cdots\hskip 85.35826pt\forall a\in\mathcal{C}$ (10)
where the ellipsis denotes the sum of other Cardy boundary states present in the theory. If there is a defect line with a non-integer quantum dimension, it turns out that one can not find any strongly symmetric boundary state Choi:2023PRD; Heymann:2024vvf.
Cloaking boundary.
Apart from the standard Cardy boundary conditions, we also consider a special class of boundary states known as cloaking boundary states Donnelly:2018ppr; Hung:2019bnq; Brehm:2022JPA; Brehm:2024arXiv. These boundary conditions play a central role in symmetry resolution because they are constructed to be invisible to a chosen set of topological defect lines. Let $\mathcal{J}\subset\mathcal{C}$ denote a closed sub-fusion category generated by a set of simple objects (which may include both invertible and non-invertible defect lines). We also define $\mathcal{I}\subset\mathrm{Irr}(\mathcal{C})$ as the collection of all simple objects whose Ishibashi states survive the projection onto the $\mathcal{J}$-subsector; equivalently, $\mathcal{I}$ consists of all labels appearing in the fusion closure of $\mathcal{J}$.
A boundary condition is said to be cloaked with respect to $\mathcal{J}$ if its Ishibashi components are invisible to the fusion action of every line in $\mathcal{J}$, i.e., the defect line passes through the boundary, unlike the Cardy states, where the lines end. This invisibility condition is imposed by requiring that the row of the modular $S$-matrix indexed by $x$ agrees with the vacuum row when restricted to the columns labeled by $\mathcal{J}$. Explicitly, the allowed Ishibashi labels $x\in\mathcal{I}$ satisfy
$\left\{x\in\mathcal{I},\qquad\frac{S_{xj}}{S_{x1}}=\frac{S_{1j}}{S_{11}}\ \ \forall\,j\in\mathcal{J}\right\}.$ (2.8)
[…]
$\bigl\|\gamma(\delta)\bigr\rangle\!\bigr\rangle=\delta_{0}\,\mathrm{Dim}(\mathcal{J})\sum_{x\in\mathcal{I}}\sqrt{S_{x1}}\,|x\rangle\!\rangle,$ (11)
where the total quantum dimension of the sub–fusion category $\mathcal{J}$ is defined by
$\mathrm{Dim}(\mathcal{J})=\sqrt{\sum_{a\in\mathcal{J}}d_{a}^{\,2}}\,.$
In the following section, we explain how cloaking boundary states are incorporated into the computation of symmetry-resolved entanglement entropy and discuss the role played by different choices of $\mathcal{J}$.
3 Symmetry resolved entanglement entropy (SREE)
Figure 2: For a set of intervals A and B, as the Hilbert space is not a direct product, we have put conformal boundary conditions with two small $\epsilon$ circles at the entangling surface, which turn into two different Cardy boundary states after conformal mapping to an annulus.
We consider a $(1+1)$ dimensional quantum system endowed with an internal symmetry and partition it into two complementary subsystems, $A$ and $B$. In order to compute the entanglement entropy of subsystem A of length $L$ with B, we have traced out the subsystem B, and this configuration is conformally equivalent to an annulus of width $W=2\log(\frac{L}{\epsilon})$, see Fig. (2). Let $\mathcal{Q}$ denote the conserved charge operator generating the symmetry. We assume that the total charge decomposes additively as
$\mathcal{Q}=\mathcal{Q}_{A}\oplus\mathcal{Q}_{B}.$ (12)
If the global state of the system is an eigenstate of $\mathcal{Q}$, it follows that
$\rho,\mathcal{Q}]=0,$ (13)
where $\rho$ is the density matrix of the full system. By tracing over the degrees of freedom of $B$, we obtain the reduced density matrix $\rho_{A}=\mathrm{Tr}_{B}\,\rho$, which commutes with the subsystem charge operator $\mathcal{Q}_{A}$. Consequently, $\rho_{A}$ has a block-diagonal structure (where each block is denoted by $\mathcal{R}_{\rho}$), with each block labeled by an eigenvalue $q_{\rho}$} of $\mathcal{Q}_{A}$:
$\rho_{A}=\bigoplus_{i}\rho_{A}(q_{\rho})\,.$ (14)
Each block $\rho_{A}(q_{\rho})$ is obtained by projecting $\rho_{A}$ onto the eigenspace of $\mathcal{Q}_{A}$ with corresponding eigenvalue. Through this decomposition, one can define symmetry-resolved entanglement measures. Following Goldstein:2017bua; german3; Bonsignori:2019naz; DiGiulio:2022jjd, the symmetry-resolved Rényi entropy, which quantifies the entanglement between subsystems $A$ and $B$ within a fixed charge sector $\mathcal{R}_{\rho}$, is defined as
$S_{n}(\mathcal{R}_{\rho}):=\frac{1}{1-n}\log\mathrm{Tr}\,\rho_{A}(q_{\rho})^{n}=\frac{1}{1-n}\log\!\left(\frac{Z(q^{n},\mathcal{R}_{\rho})}{Z^{n}(q,\mathcal{R}_{\rho})}\right).$ (15)
Here, $Z(q)$ denotes the full annulus partition function in the open-string channel and closed-string channel related by modular $S$ transformation Cardy:2004hm; Petkova:2000ip, expressed as  [^7] In light of Brehm:2024arXiv, and as we will also observe in later computations, the open–closed duality is anomalous: the correspondence does not hold exactly for the cloaking boundary state. This anomaly appears in the open-string spectrum, where some coefficients become non-integer. Nevertheless, in our analysis we continue to use this duality, and the final result still admits a consistent probabilistic interpretation. Strictly speaking, the anomaly may introduce an additional correction to the SREE, but this can be absorbed into the universal term. We thank Brandon Rayhaun for emphasizing this point. ,
[…]
where $B_{1}$ and $B_{2}$ label the boundary Cardy states and $q=e^{-\frac{2\pi^{2}}{W}},\tilde{q}=e^{-2W}$. Throughout this work, we adopt the notation $Z(q)$ in place of $Z_{a|b}(q)$. In subsequent sections, the structure and interpretation of the coefficients $n_{ia}^{b}$ will be elaborated in detail. Accordingly, $Z(q,\mathcal{R}_{\rho})$ refers to the sector of the full partition function corresponding to the symmetry representation denoted by $\mathcal{R}_{\rho}$. For 1+11+1D RCFT, the global symmetry is generated by topological defect lines (TDLs), and they form a fusion category. In order to resolve the entanglement Rényi entropy with respect to the fusion category of the underlying TDLs, the partition functions of each irreducible sectors can be extracted from the full annulus partition function by looking at the action of the TDLs on the primary fields rather than computing charged moments DiGiulio:2022jjd [^8] For the computation of charged moments, we have to compute the projectors in terms of the $F$-symbols. For generic RCFT, it is very hard to compute these symbols. For more detail about the $F$-symbols upto rank $7$, see Ardonne_2010; vercleyen2023lowrankfusionrings..
Boundary conditions and SREE
In this section, we briefly review some subtleties in the choice of boundary states and their consequences for the symmetry-resolved Rényi entropies (SRRE). For illustration, suppose there are $N$ irreducible blocks; the Rényi entropy for the $\rho$’th block admits the following asymptotic behaviour in the high energy limit $q\to 1$ or $\tilde{q}\to 0$,
$\begin{split}S_{n}(\mathcal{R}_{\rho})&=\frac{1}{1-n}\log\!\left(\frac{Z(q^{n},\mathcal{R}_{\rho})}{Z^{n}(q,\mathcal{R}_{\rho})}\right),\qquad\frac{Z(q^{n},\mathcal{R}_{\rho})}{Z^{n}(q,\mathcal{R}_{\rho})}\xrightarrow{\tilde{q}\to 0}p_{\rho}^{\,n-1}\,\tilde{q}^{\big(n-\frac{1}{n}\big)\frac{c_{\rm eff}}{24}},\end{split}$ (17)
with
$Z(q^{n},\mathcal{R}_{\rho})\overset{\tilde{q}\to 0}{\sim}p_{\rho}\,\chi_{\textrm{GS}}(\tilde{q}^{1/n})$ (18)
where $c_{\rm eff}=c-24\Delta_{\rm GS}$ and $\Delta_{\rm GS}$ is the conformal dimension of the ground state (for unitary theory ground state is the vacuum state with $\Delta_{\rm GS}=0$, but non-zero for non-unitary theory). In this limit, the SRRE for the irreducible block becomes
$S_{n}(\mathcal{R}_{\rho})=\frac{c_{\rm eff}}{6}\left(\frac{n+1}{n}\right)\log\!\left(\frac{L}{\epsilon}\right)+\log\!\big(p_{\rho}\big)+\cdots.$ (19)
From this we can get the symmetry-resolved entanglement entropy (SREE) by taking $n\rightarrow 1$ limit. Our aim is not only to compute the SREE but also to relate the $p_{\rho}$’s to modular data and topological quantities. We therefore propose the following prescriptions for different classes of boundary states.
Weakly symmetric boundary condition
For a weakly symmetric boundary state, the SREE in the charge sector labelled by $\rho\in\texttt{Irr}(\mathcal{C})$ takes the form
${S}(\mathcal{R}_{\rho})$ $=\frac{c_{\rm eff}}{3}\log\!\left(\frac{L}{\epsilon}\right)+\log(g_{1}g_{2})+\log\!\left(\frac{d_{\rho}\,N_{\rho B_{1}}^{B_{2}}}{d(B_{1}\otimes B_{2})}\right),$ (20)
where the $g_{i}$’s are boundary entropies and, if a block contains several irreducible elements (this scenario may appear when the full resolution is not possible), a sum over the block index $\rho$ is implied. For multiplicity-free fusion of the defect lines, one finds
$d(B_{1}\otimes B_{2})=\sum_{\rho}d_{\mathcal{L}_{\rho}},$ (21)
so that (20) reduces to
${S}(\mathcal{R}_{\rho})=\frac{c_{\rm eff}}{3}\log\!\left(\frac{L}{\epsilon}\right)+\log(g_{1}g_{2})+\log\!\left(\frac{d_{\rho}}{\sum_{\rho\in\texttt{Irr}(\mathcal{C})}d_{\rho}}\right).$ (22)
Remark. For a weakly symmetric boundary, one may regard $\rho$ as an element of $B_{1}\otimes B_{2}$. The fusion product $B_{1}\otimes B_{2}$ then generates the entire subfusion sector. Equation (22) holds essentially in multiplicity-free situations, i.e. when there are no fusion multiplicities and each irreducible block carries a single character.
The expression (20) can be viewed as a special instance of a more general relation,
$S(\mathcal{R}_{\rho})=\frac{c_{\rm eff}}{3}\log\!\left(\frac{L}{\epsilon}\right)+\log(g_{1}g_{2})+\log\!\left(\frac{d_{\rho}\,\overline{\mathcal{N}}_{\rho B_{1}}^{\,\,\,B_{2}}}{d_{B_{1}}d_{B_{2}}}\right),$ (23)
which applies to a broader class of boundary conditions. Here $\overline{\mathcal{N}}_{\rho B_{1}}^{\,\,B_{2}}$ denotes the boundary fusion coefficient defined via the action of a topological defect line on a boundary state:
$\mathcal{L}_{\rho}\,\|B_{1}\rangle\!\rangle=\overline{\mathcal{N}}_{\rho B_{1}}^{\,\,B_{2}}\,\|B_{2}\rangle\!\rangle+\sum|\!|CS\rangle\!\rangle.$ (24)
Using the state–operator correspondence together with the Cardy construction, one finds the explicit action (for diagonal RCFT) Petkova:2000ip,
$\mathcal{L}_{\rho}\|B_{i}\rangle\!\rangle=\sum_{j=1}^{N}\frac{S_{ij}}{\sqrt{S_{1j}}}\,\mathcal{L}_{\rho}|j\rangle\!\rangle=\sum_{j=1}^{N}\frac{S_{ij}}{\sqrt{S_{1j}}}\,\frac{S_{\rho j}}{S_{1j}}\,|j\rangle\!\rangle\,.$ (25)
Note that both the Cardy and Ishibashi states are not primary states. The Ishibashi state can be expanded as a sum over a primary and its descendants:
$|\mu\rangle\!\rangle=\sum_{N=0}^{\infty}|\mu,N\rangle\otimes|\bar{\mu},N\rangle,$ (26)
where $\{|\mu,N\rangle\}$ is an orthonormal basis of the irreducible module $\mathcal{H}_{\mu}$. The primary state is $|\mu,0\rangle$, and the Virasoro descendants are generated as
$|\mu,N\rangle=L_{-n_{1}}L_{-n_{2}}\cdots L_{-n_{k}}\,|\mu,0\rangle,\qquad n_{1}+n_{2}+\cdots+n_{k}=N\,.$ (27)
Consider now the action of $\mathcal{L}_{\rho}$ on the Ishibashi state:
$\begin{split}\mathcal{L}_{\rho}|\mu\rangle\!\rangle&=\sum_{N=0}^{\infty}\mathcal{L}_{\rho}|\mu,N\rangle\otimes|\bar{\mu},N\rangle=\sum_{N=0}^{\infty}\mathcal{L}_{\rho}\,L_{-n_{1}}L_{-n_{2}}\cdots L_{-n_{k}}|\mu,0\rangle\otimes|\bar{\mu},N\rangle\,,\\ &=\sum_{N=0}^{\infty}L_{-n_{1}}L_{-n_{2}}\cdots L_{-n_{k}}\mathcal{L}_{\rho}|\mu,0\rangle\otimes|\bar{\mu},N\rangle\,,\\ &\overset{\texttt{state-operator}}{=}\sum_{N=0}^{\infty}L_{-n_{1}}L_{-n_{2}}\cdots L_{-n_{k}}\underbrace{\mathcal{L}_{\rho}\Phi_{\mu}}_{S_{\rho\mu}/S_{1\mu}}|0\rangle\otimes|\bar{\mu},N\rangle\longrightarrow\frac{S_{\rho\mu}}{S_{1\mu}}\,|\mu\rangle\!\rangle\,.\end{split}$ (28)
Since $\mathcal{L}_{\rho}$ is a symmetry generator, it commutes with all Virasoro generators. Thus, even though Ishibashi states are not primary, the action of $\mathcal{L}_{\rho}$ on those states mimics that of primary fields, unlike Cardy states. From (25), one can then recombine terms to identify the coefficients $\overline{\mathcal{N}}_{\rho B_{1}}^{B_{2}}$. Open–closed duality further imposes the consistency condition
$d_{B_{1}}d_{B_{2}}=\sum_{\rho\in\texttt{Irr}(\mathcal{C})}\overline{\mathcal{N}}_{\rho B_{1}}^{\,\,\,B_{2}}\,d_{\rho}\,.$ (29)
Verification of (29) for real, symmetric modular data
The annulus partition function in the open-string channel is
[…]
In the closed channel, the corresponding partition function reads
[…]
Matching the leading terms from both channels yields
$\frac{1}{S_{11}}\sum_{\rho}\overline{\mathcal{N}}_{\rho B_{1}}^{\,\,B_{2}}S_{\rho 1}=d_{B_{1}}d_{B_{2}}\quad\Longrightarrow\quad d_{B_{1}}d_{B_{2}}=\sum_{\rho}\overline{\mathcal{N}}_{\rho B_{1}}^{\,\,B_{2}}d_{\rho},$ (32)
as claimed in (29). A similar derivation holds for more general modular data with complex entries.
Why (20) holds for weakly symmetric boundaries
For weakly symmetric Cardy boundaries, the action of topological defect lines on boundary Ishibashi states closes according to the bulk fusion algebra:
$\mathcal{L}_{\rho}\,\|\Phi_{i}\rangle\!\rangle=\sum_{j=1}^{N}N_{\rho i}^{\;j}\,\|\Phi_{j}\rangle\!\rangle,$ (33)
where the bulk fusion coefficients are given by
$\Phi_{\rho}\times\Phi_{i}=\sum_{j=1}^{N}N_{\rho i}^{\;j}\,\Phi_{j},\qquad N_{\rho i}^{\;j}=\sum_{n=1}^{N}\frac{S_{\rho n}\,S_{in}\,S_{nj}^{*}}{S_{1n}}.$ (34)
Although there is no strict one-to-one correspondence between Cardy boundary states and bulk primaries, the structural similarity between (33) and (34) justifies identifying the defect action with the associated fusion data. This correspondence explains the validity of the simpler formula (20) in the weakly symmetric case.
Strongly (and possibly cloaking) boundary condition
The prescription (22) fails in the presence of strongly symmetric boundary states. For a strongly symmetric boundary state, the topological defect line acts simply by its quantum dimension:
$\mathcal{L}_{\rho}\,\|B\rangle\!\rangle=d_{\rho}\,\|B\rangle\!\rangle\quad\Longrightarrow\quad\overline{\mathcal{N}}_{\rho B}^{\;B}=d_{\rho}\,.$ (35)
Substituting this relation into (23) and using the open–closed identity (29) yields the SREE  [^9] For cloaking states the formula in (36) matches with Sierra_2024; Choi:2024wfm.
$S(\mathcal{R}_{\rho})=\frac{c_{\rm eff}}{3}\log\!\left(\frac{L}{\epsilon}\right)+\log(g_{1}g_{2})+\log\!\left(\frac{d_{\rho}^{2}}{\sum_{\rho\in\texttt{Irr}(\mathcal{C})}d_{\rho}^{2}}\right)\,.$ (36)
The formula (36) is expected to apply also to the cloaking boundary state Brehm:2022JPA; Brehm:2024arXiv introduced in (11). In many diagonal RCFTs, the cloaking state coincides with the identity Ishibashi state $|\mathbb{I}\rangle\!\rangle$ Brehm:2022JPA; Brehm:2024arXiv, for which the defect acts by its quantum dimension:
$\mathcal{L}_{\rho}\,|\mathbb{I}\rangle\!\rangle=d_{\rho}\,|\mathbb{I}\rangle\!\rangle,$ (37)
so that $\overline{\mathcal{N}}_{\rho\,\mathbb{I}}^{\,\mathbb{I}}=d_{\rho}$ and Eq. (36) follows directly. We stress, however, that although (36) holds in these (cloaking) examples, the derivations (e.g (32)) employed for the Cardy boundary states do not carry over in a straightforward way to cloaking boundaries (as cloaking states are not physical Cardy states).
4 Computing SREE in Unitary Rational CFT
In this section, we investigate symmetry resolution in both unitary diagonal and non-diagonal RCFTs. For diagonal models, we primarily focus on the Tambara-Yamagami (TY) symmetry resolution in the Ising and tetracritical Ising CFTs. For non-diagonal RCFTs, we examine two representative cases: the $E_{6}$ theory and its $\tfrac{1}{2}E_{6}$ fusion ring, and the $\mathbb{Z}_{3}$ symmetry resolution in the three-state Potts model.
4.1 Symmetry resolution with respect to Tambara-Yamagami (TY) Fusion Category for Ising CFT
The simplest unitary Virasoro minimal model, namely the critical Ising model, is the prototype example for two-dimensional rational CFT. It has three primaries, $\{\mathbb{I},\varepsilon,\sigma\}$, and their non-trivial fusions are
$\varepsilon\times\varepsilon=\mathbb{I},\quad\varepsilon\times\sigma=\sigma,\quad\sigma\times\sigma=\mathbb{I}+\varepsilon\,.$ (38)
This is a diagonal rational CFT. So, it has three topological defects labeled by the primaries, namely, $\{\mathcal{L}_{\mathbb{I}},\mathcal{L}_{\mathbb{\varepsilon}},\mathcal{L}_{\mathbb{\sigma}}\}=\mathcal{C}$ and their action on the primaries is summarized in Table (1).
Defect line $\mathbb{I}$ $\varepsilon$ $\sigma$
$\mathcal{L}_{\mathbb{I}}$ 11 11 11
$\mathcal{L}_{\mathbb{\varepsilon}}$ 11 11 1-1
$\mathcal{L}_{\mathbb{\sigma}}$ $\sqrt{2}$ $\sqrt{2}$ 0
Table 1: Action of the TY category on the primaries .
We want to study symmetry-resolved entanglement entropy for this TY fusion category symmetry. In order to do that, we need a strongly symmetric or weakly symmetric entangled boundary condition according to Choi:2023PRD; Choi:2024wfm; Heymann:2024vvf; Das:2024qdx. But it turns out that for this case, there is no such boundary at the entangled surfaces. However, one can use the cloaking boundary condition Brehm:2022JPA; Brehm:2024arXiv for the purpose of symmetry resolution. For the critical Ising model, it turns out that the Ishibashi state gives the cloaking boundary state that corresponds to the identity primary Brehm:2024arXiv
$|\!|\gamma(\delta)\rangle\!\rangle=|\mathbb{I}\rangle\!\rangle\,.$ (39)
Now, we can compute the annulus partition function in the closed string channel when we sandwich the cloaking state (39) at the boundary.
[…]
The TY fusion category has three elements, and there are three representations (each irreducible block is represented by $\mathcal{R}_{\rho},\,\rho=1,2,3$). Hence, following the discussion presented in Sec. (3), the reduced density matrix can be decomposed into three blocks. From the Table (1), we can see the action of $\mathcal{C}$ on each primary is different; therefore, we can write the symmetry-resolved partition functions as,
$\begin{split}Z(q^{n},\mathcal{R}_{1})=\frac{\frac{1}{2}\chi_{1}(q^{n})}{Z^{n}(q)}\,,\quad Z(q^{n},\mathcal{R}_{2})=\frac{\frac{1}{2}\chi_{2}(q^{n})}{Z^{n}(q)}\,,\quad Z(q^{n},\mathcal{R}_{3})=\frac{\frac{1}{\sqrt{2}}\chi_{3}(q^{n})}{Z^{n}(q)}\,.\end{split}$ (40)
Now, using (15), we can compute the symmetry-resolved $n$th Rényi . As discussed in (17), in the high-energy limit, the leading contribution comes from the vacuum. Therefore,
$\frac{Z(q^{n},\mathcal{R}_{1})}{Z^{n}(q,\mathcal{R}_{1})}\overset{\tilde{q}\to 0}{\sim}\left(\frac{1}{4}\right)^{1-n}\widetilde{q}^{\,(n-\frac{1}{n})\frac{c}{24}}$ (41)
and SREE for this block becomes
$S(\mathcal{R}_{1})=\frac{c}{6}\left(\frac{n+1}{n}\right)\log\left(\frac{L}{\epsilon}\right)+\log\frac{1}{4}+\ldots$ (42)
where the constant term
$\log\frac{1}{4}=\log\left(\frac{d_{\mathcal{R}_{1}}^{\,2}}{\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}^{\,2}}\right)+\log(\langle\langle\gamma(\delta)||\Omega\rangle\langle\Omega|\gamma(\delta)\rangle\rangle)\,.$ (43)
where, $d_{\mathcal{R}_{1}}=1$, $\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}^{2}=4$, and $\log(\langle\!\langle\gamma(\delta)|\!|\Omega\rangle\!\langle\Omega|\!|\gamma(\delta)\rangle\!\rangle)=0$ corresponds to the Affleck-Ludwig boundary entropy Affleck:1991tk. Similar result can be obtained for $S(\mathcal{R}_{2})$, and
$S(\mathcal{R}_{3})=\frac{c}{6}\left(\frac{n+1}{n}\right)\log\left(\frac{L}{\epsilon}\right)+\log\frac{1}{2}+\ldots\,$, (44)
where the constant term
$\log\frac{1}{2}=\log\left(\frac{d^{\,2}_{\mathcal{R}_{3}}}{\sum\limits_{\mathcal{R}_{i}}d^{\,2}_{\mathcal{R}_{i}}}\right)+\log(\langle\!\langle\gamma(\delta)|\!|\Omega\rangle\langle\Omega|\!|\gamma(\delta)\rangle\!\rangle).$ (45)
Here $d_{\mathcal{R}_{3}}=\sqrt{2}$ is the quantum dimension of $\mathcal{L}_{\sigma}$. We observe that equation (36) is satisfied in this example when using the specified boundary states, known as the cloaking states. Through a detailed examination of this explicit case, it becomes clear that symmetry resolution in the Ising CFT with respect to the TY fusion ring cannot be achieved using the standard procedure, as no weakly symmetric state exists in this case. However, by employing the cloaking state, we demonstrate that such a resolution becomes feasible. To the best of our knowledge, this work presents the first TY-resolution of the Ising CFT in the literature.
4.2 Symmetry resolution for $\mathcal{M}(6,5)$
We now analyze the symmetry resolution of the diagonal $\mathcal{M}(6,5)$ theory. Interestingly, in this case, there are two natural choices of sub-fusion categories, which we treat separately. For the larger sub-fusion category we find that no weakly symmetric boundary state exists, but we can still perform the resolution using an appropriate cloaking boundary state. In contrast, for the smaller sub-fusion category, a weakly symmetric boundary state is available, and the symmetry resolution proceeds accordingly.
Tetracritical Ising model. The diagonal minimal model corresponding to the modular data of $\mathcal{M}(6,5)$, the tetracritical Ising model Chang:2018iay; Choi:2023PRD, is another example of a diagonal RCFT. The model $\mathcal{M}(6,5)$ has central charge $c=\tfrac{4}{5}$ and admits ten simple Verlinde lines, which we label as  [^10] In contrast, for the Potts CFT, one of the simple TDLs is $N$, satisfying $N\times N=1+\eta+\bar{\eta}$, which is a $\mathbb{Z}_{3}$-TY line. In the diagonal theory, however, we instead encounter the composite line $M=\eta+\bar{\eta}$ with $N\times N=I+M$, which is not strictly of TY type.
${I},\;C=\mathcal{L}_{Y},\;M=\mathcal{L}_{Z},\;W=\mathcal{L}_{X},\;MW,\;CW,\;N,\;CN,\;WN,\;CWN.$ (46)
Among these lines the subset $\{\mathcal{L}_{\mathbb{I}},\,C,\,N,\,CN,\,M\}$ closes into a TY-type sub-fusion with the nontrivial fusion rules
$C^{2}=I,\qquad M^{2}=I+M+C,\qquad N^{2}=I+M.$ (47)
We study symmetry resolution with respect to these five simple TDLs. The full fusion category contains ten bulk primaries,
$\mathbb{I},\;Y,\;\varepsilon,\;X,\;Z,\;\sigma,\;\Phi_{4,4},\;\Phi_{2,2},\;\Phi_{4,2},\;\Phi_{3,2}.$ (48)
To identify the irreducible charge blocks, one first determines the action of the TDLs on these primaries; this action is displayed in the Table (2).
Defect line $\mathbb{I}$ $Y$ $\varepsilon$ $X$ $Z$ $\sigma$ $\Phi_{4,4}$ $\Phi_{2,2}$ $\Phi_{4,2}$ $\Phi_{3,2}$
II 11 11 11 11 11 11 11 11 11 11
$M$ 22 22 22 22 1-1 1-1 0 0 0 0
$C$ 11 11 11 11 11 11 1-1 1-1 1-1 1-1
NN $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ 0 0 1-1 1-1 11 11
CNCN $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ $\sqrt{3}$ 0 0 11 11 1-1 1-1
Table 2: Action of the TDLs on the $\mathcal{M}(6,5)$ primaries of the tetracritical Ising model.
From Table (2), one sees that there are five irreducible blocks, each containing two primaries. In the absence of symmetric (weakly symmetric) boundary states, we choose the cloaking boundary, which we take to be the identity Ishibashi state, $\|\gamma(\delta)\rangle\!\rangle=\|\mathbb{I}\rangle\!\rangle.$ The annulus partition function in the closed-string channel is
$\begin{split}Z(\tilde{q})&=\langle\!\langle\gamma(\delta)\|\tilde{q}^{L_{0}-\frac{c}{24}}\|\gamma(\delta)\rangle\!\rangle=\chi_{1}(\tilde{q})\,,\\ &\overset{!}{=}\sqrt{\frac{1}{15}\!\left(\frac{5}{8}-\frac{\sqrt{5}}{8}\right)}\chi_{\mathbb{I}}(q)+\sqrt{\frac{1}{15}\!\left(\frac{5}{8}-\frac{\sqrt{5}}{8}\right)}\chi_{Y}(q)+\sqrt{\frac{1}{15}\!\left(\frac{\sqrt{5}}{8}+\frac{5}{8}\right)}\chi_{\varepsilon}(q)\\ &\quad+\sqrt{\frac{1}{15}\!\left(\frac{\sqrt{5}}{8}+\frac{5}{8}\right)}\chi_{X}(q)+2\sqrt{\frac{1}{15}\!\left(\frac{5}{8}-\frac{\sqrt{5}}{8}\right)}\chi_{Z}(q)+2\sqrt{\frac{1}{15}\!\left(\frac{\sqrt{5}}{8}+\frac{5}{8}\right)}\chi_{\sigma}(q)\\ &\quad+\sqrt{\frac{1}{5}\!\left(\frac{5}{8}-\frac{\sqrt{5}}{8}\right)}\chi_{4,4}(q)+\sqrt{\frac{1}{5}\!\left(\frac{5}{8}-\frac{\sqrt{5}}{8}\right)}\chi_{4,2}(q)\\ &\quad+\sqrt{\frac{1}{5}\!\left(\frac{\sqrt{5}}{8}+\frac{5}{8}\right)}\chi_{2,2}(q)+\sqrt{\frac{1}{5}\!\left(\frac{\sqrt{5}}{8}+\frac{5}{8}\right)}\chi_{3,2}(q)\,.\end{split}$ (49)
The leading contribution to the block partition functions in the $\tilde{q}\to 0$ limit is
$\begin{split}&Z(q^{n},\mathcal{R}_{1})\sim\tfrac{1}{12}\,\chi_{\mathbb{I}}(\tilde{q}^{1/n}),\qquad Z(q^{n},\mathcal{R}_{2})\sim\tfrac{1}{12}\,\chi_{\mathbb{I}}(\tilde{q}^{1/n}),\\ &Z(q^{n},\mathcal{R}_{3})\sim\tfrac{1}{3}\,\chi_{\mathbb{I}}(\tilde{q}^{1/n}),\qquad Z(q^{n},\mathcal{R}_{4})\sim\tfrac{1}{4}\,\chi_{\mathbb{I}}(\tilde{q}^{1/n}),\qquad Z(q^{n},\mathcal{R}_{5})\sim\tfrac{1}{4}\,\chi_{\mathbb{I}}(\tilde{q}^{1/n})\,.\end{split}$ (50)
Therefore, the symmetry (Cat-)SREE for the $\rho$-th irreducible block is
$\begin{split}S_{n}(\mathcal{R}_{\rho})&=\frac{1}{1-n}\log\!\left(\frac{Z(q^{n},\mathcal{R}_{\rho})}{Z^{n}(q,\mathcal{R}_{\rho})}\right),\quad\text{with}\quad\frac{Z(q^{n},\mathcal{R}_{\rho})}{Z^{n}(q,\mathcal{R}_{\rho})}\xrightarrow{\tilde{q}\to 0}p_{\rho}^{\,n-1}\,\tilde{q}^{\big(n-\frac{1}{n}\big)\frac{c_{\rm eff}}{24}},\end{split}$ (51)
where the coefficients are $p_{\rho}=\tfrac{1}{12},\tfrac{1}{12},\tfrac{1}{3},\tfrac{1}{4},\tfrac{1}{4}$ for the five irreducible blocks, respectively. In the cloaking (identity Ishibashi) boundary choice, this yields the compact form as (36),
${S}_{n}(\mathcal{R}_{\rho})=\frac{c_{\rm eff}}{6}\left(\frac{n+1}{n}\right)\log\!\left(\frac{L}{\epsilon}\right)+\log(g^{2})+\log\!\left(\frac{d_{\rho}^{2}}{\sum_{\rho}d_{\rho}^{2}}\right),\,\textrm{with,}\,\,g=1$ (52)
with the quantum dimensions for the five lines given by
$d_{\mathbb{I}}=1,\quad d_{M}=2,\quad d_{C}=1,\quad d_{N}=\sqrt{3},\quad d_{CN}=\sqrt{3}.$
Hence
$\sum_{\rho}d_{\rho}^{2}=1^{2}+2^{2}+1^{2}+(\sqrt{3})^{2}+(\sqrt{3})^{2}=12,$ (53)
which reproduces the numerical prefactors $p_{\rho}=\tfrac{1}{12},\tfrac{1}{12},\tfrac{1}{3},\tfrac{1}{4},\tfrac{1}{4}$ quoted above for the five irreducible blocks.
Resolution with weakly symmetric boundary state:
As observed above, symmetric boundary states do not exist within the sub-fusion sector under consideration. However, one can identify a weakly symmetric boundary state by resolving the minimal model $\mathcal{M}(6,5)$ with respect to the smaller subset of Verlinde lines $\{I,\,C,\,M\}$. From the fusion algebra restricted to this subset, it is immediate that the boundary state $\|M\rangle\!\rangle$ satisfies the weak symmetry condition. Consequently, the annulus partition function becomes
$Z(\tilde{q})=\langle\!\langle M\|\;\tilde{q}^{L_{0}-\frac{c}{24}}\;\|M\rangle\!\rangle=\chi_{1}(q)+\chi_{Y}(q)+\chi_{Z}(q).$ (54)
The action of the topological defect lines shows that there are three irreducible blocks. Their annulus block partition functions in the $\tilde{q}\to 0$ limit behave as
$\begin{split}Z(q^{n},\mathcal{R}_{1})&\sim\tfrac{1}{2}\sqrt{\tfrac{1}{30}\!\left(5-\sqrt{5}\right)}\;\chi_{1}(\tilde{q}^{1/n}),\,\,\,\,\,\,Z(q^{n},\mathcal{R}_{2})\sim\tfrac{1}{2}\sqrt{\tfrac{1}{30}\!\left(5-\sqrt{5}\right)}\;\chi_{1}(\tilde{q}^{1/n}),\\ &\hskip 56.9055ptZ(q^{n},\mathcal{R}_{3})\sim\sqrt{\tfrac{1}{30}\!\left(5-\sqrt{5}\right)}\;\chi_{1}(\tilde{q}^{1/n}).\end{split}$ (55)
The boundary entropy for this choice is
$\log(g^{2})=\log|\langle\!\langle M\|\,\mathbb{I}\rangle\!\rangle\big|^{2}=\log\!\sqrt{\tfrac{2}{15}\!\left(5-\sqrt{5}\right)}.$ (56)
With the weakly symmetric boundary $\|M\rangle\!\rangle$ the Cat-SREE in block $\rho$ reads
$\begin{split}S_{n}(\mathcal{R}_{\rho})&=\frac{c_{\rm eff}}{6}\left(\frac{n+1}{n}\right)\log\!\left(\frac{L}{\epsilon}\right)+\log(g^{2})+\log\!\left(\frac{d_{\rho}\,N_{\rho B_{1}}^{B_{2}}}{d(B_{1}\otimes B_{2})}\right),\end{split}$ (57)
where, for the lines under consideration,
$d_{\mathbb{I}}=1,\quad d_{C}=1,\quad d_{M}=2,$ (58)
and the relevant fusion coefficients are
$N_{\mathbb{I},Z}^{\,Z}=1,\qquad N_{Y,Y}^{\,Z}=1,\qquad N_{Z,Z}^{\,Z}=1.$ (59)
Therefore, the symmetry (Cat-)SREE for the $\rho$-th irreducible block is
$\begin{split}S_{n}(\mathcal{R}_{\rho})&=\frac{1}{1-n}\log\!\left(\frac{Z(q^{n},\mathcal{R}_{\rho})}{Z^{n}(q,\mathcal{R}_{\rho})}\right),\quad\text{with}\quad\frac{Z(q^{n},\mathcal{R}_{\rho})}{Z^{n}(q,\mathcal{R}_{\rho})}\xrightarrow{\tilde{q}\to 0}p_{\rho}^{\,n-1}\,\tilde{q}^{\big(n-\frac{1}{n}\big)\frac{c_{\rm eff}}{24}},\end{split}$ (60)
where the coefficients are $p_{\rho}=\tfrac{1}{6},\tfrac{1}{6},\tfrac{1}{3}$ for the five irreducible blocks respectively.
4.3 Symmetry resolution in non-diagonal RCFTs
Before turning to non-unitary RCFTs, we take a pause and examine key features of symmetry resolution for non-invertible symmetries in non-diagonal rational conformal field theories. Through an explicit example, we highlight the conceptual and technical challenges that distinguish non-diagonal theories from their diagonal counterparts. In particular, we illustrate how symmetry resolution may fail in non-diagonal RCFTs by studying the non-diagonal $E_{6}$ model and its associated $\tfrac{1}{2}E_{6}$ fusion ring. We also give the example of $\mathbb{Z}_{3}$ resolution in the three-state Potts model Haghighat:2023sax and show that, for this case, the resolution is possible.
$E_{6}$ and its $\tfrac{1}{2}E_{6}$ symmetry resolution?
We consider $SU(2)_{10}$ WZW model of $E_{6}$-type DiFrancesco:1997nk. The torus partition function is given by,
$Z(\tau)=|\chi_{0}+\chi_{3}|^{2}+|\chi_{\frac{3}{2}}+\chi_{\frac{7}{2}}|^{2}+|\chi_{2}+\chi_{5}|^{2}$ (61)
where, $\chi_{j}(\tau)$ is spin $j$ $SU(2)$ affine character. The modular S-matrix for $SU(2)_{10}$ is a $11\times 11$ matrix which can be constructed by the following,
$S_{jj^{\prime}}=\sqrt{\frac{2}{k+2}}\sin\left(\frac{(2j+1)(2j^{\prime}+1)\pi}{k+2}\right),\,\textrm{with,}\,\,k=10,\,\,\textrm{and,}\,j,j^{\prime}=0,\frac{1}{2},\cdots,5\,.$ (62)
The $\frac{1}{2}E_{6}$ fusion ring consists of three simple TDLs: $\mathcal{L}_{1}=I,\,\mathcal{L}_{5}=X,\,\mathcal{L}_{3}=Y$. The fusion of the TDLs can be derived from the matrix ($\psi$) constructed from the eigenvector of the adjacency matrix of $E_{6}$, which can be derived from the Dynkin diagram.
[…]
The $\psi$-matrix is given by [^11] The $a$ index in the subscript of $\psi$ should not be confused with the element $a$ of the matrix. ,
$\psi_{a}^{j}=\left(\begin{array}[]{cccccc}a&\frac{1}{2}&b&b&\frac{1}{2}&a\\ b&\frac{1}{2}&a&-a&-\frac{1}{2}&-b\\ c&0&-d&-d&0&c\\ b&-\frac{1}{2}&a&-a&\frac{1}{2}&-b\\ a&-\frac{1}{2}&b&b&-\frac{1}{2}&a\\ d&0&-c&c&0&-d\\ \end{array}\right)$ (70)
where, $a=\frac{1}{2}\sqrt{\frac{1}{6}\left(3-\sqrt{3}\right)},\,b=\frac{1}{2}\sqrt{\frac{1}{6}\left(\sqrt{3}+3\right)},\,c=\frac{1}{2}\sqrt{\frac{1}{3}\left(\sqrt{3}+3\right)},\,d=\frac{1}{2}\sqrt{\frac{1}{3}\left(3-\sqrt{3}\right)}\,.$ Here, $a$ runs over the nodes of the corresponding Dynkin diagram and $j$ runs over the exponents. The fusion can be computed from the modified Verlinde formula DiFrancesco:1997nk; Behrend:1998mu; Behrend:1999bn; Petkova:2000ip,
$\hat{\mathcal{N}}_{ab}^{c}\text{:=}\sum_{j\in\textrm{Exp}(G)}\frac{\psi_{a}^{j}\,\,\psi_{b}^{j}\,\,\psi_{c}^{j}}{\psi_{1}^{j}}$ (71)
[…]
$X^{2}=I,\quad Y^{2}=I+X+2Y,\quad XY=YX=Y\,.$ (72)
[…]
where, $n_{ia}^{b}=\sum_{j\in\textrm{Exp}(G)}\frac{\psi_{a}^{j}\,\psi_{b}^{j}\,S_{ij}}{S_{1j}}$. From (73) we observe that the open-string spectrum generated by the boundary $|\!|Y\rangle\!\rangle$ contains SU(2)10 primaries which are absent from the $E_{6}$, i.e., the open string spectrum is not solely inside the physical spectrum of the theory. As a result, the $\tfrac{1}{2}E_{6}$ fusion ring does not act by scalars on those extra open-channel primaries. To be precise, $\mathcal{L}_{a}\phi_{j}=\frac{\psi_{a}^{j}}{\psi_{1}^{j}}\phi_{j},\,\mathcal{L}_{i}\phi_{j}=\frac{S_{ij}}{S_{1j}}\phi_{j}$, with, $j\in\textrm{Exp}(G),\,i\in\mathcal{I}$. Hence symmetry resolution based solely on the intrinsic $\tfrac{1}{2}E_{6}$ fusion data is not canonical: one may still decompose the open spectrum with respect to the defect algebra, but the decomposition is in general non-diagonal (action on the defects by the TDLs are via matrices rather than by scalar projectors) and requires embedding the problem into the full SU(2)10 theory. To elaborate further, $\mathcal{L}_{a}\phi_{i}=(D_{a})_{i,j}\phi_{j}$ with, $i\notin\textrm{Exp}(G)$, one can not call it a symmetry. Hence, the notion of symmetry resolution is not defined in the usual sense. As an aside, we may still compute the full entanglement entropy, which is given by,
$S=\frac{c}{3}\log\left(\frac{L}{\epsilon}\right)+\log(g^{2})+\cdots$ (74)
where the boundary entropy is given by,
$\log(g^{2})=\log|\langle\!\langle I||\!|Y\rangle\!\rangle|^{2}=\log(2+\sqrt{3})\,.$ (75)
Three state Potts and its $\mathbb{Z}_{3}$ resolution
In three state Potts model, the $\mathbb{Z}_{3}$ fusion ring consists of three simple TDLs: $I,\,\eta,\,\eta^{2}$. The fusion of these TDLs is given by,
$\eta\times\eta=\eta^{2},\,\eta\times\eta^{2}=1,\,\eta^{2}\times\eta^{2}=\eta\,.$ (76)
As we see, we do not have any weakly symmetric boundary in the fusion ring. However, we can choose a bigger fusion ring that consists of another simple object $N$, which has the following fusion Haghighat:2023sax:
$N\times N=1+\eta+\eta^{2},\,N\times\eta=N,\,N\times\bar{\eta}=N\,.$ (77)
As we see, $N$ can be treated as a strongly symmetric boundary, and the annulus open string partition function is given by Affleck:1998nq; Behrend:1998mu; Behrend:1999bn,
[…]
As we see, unlike the $E_{6}$ case, the open string spectrum is solely in the Potts spectrum. Now, we study symmetry resolution with respect to the three ($I,\,\eta,\,\eta^{2}$) invertible simple TDLs. The full fusion category contains eight bulk primaries,
II $Y$ $Z$ $Z^{*}$
II 1 1 1 1
$\eta$ 11 11 $\omega$ $\omega^{2}$
$\eta^{2}$ 11 11 $\omega^{2}$ $\omega$
Table 3: Action of TDLs on the open string spectrum primaries Sinha:2023hum; Gu:2023yhm.
From Table (3), we see that there are three irreducible blocks corresponding to the three simple TDLs: $(I,Y),\,Z,\,Z^{*}$. The leading contributions to the block partition functions in the $\tilde{q}\to 0$ limit are
$\begin{split}&Z(q^{n},\mathcal{R}_{1})\sim\sqrt{\tfrac{1}{30}(5-\sqrt{5})}\,\chi_{\mathbb{I}}(\tilde{q}^{1/n}),\quad Z(q^{n},\mathcal{R}_{2})\sim\sqrt{\tfrac{1}{30}(5-\sqrt{5})}\,\chi_{\mathbb{I}}(\tilde{q}^{1/n}),\\ &\hskip 85.35826ptZ(q^{n},\mathcal{R}_{3})\sim\sqrt{\tfrac{1}{30}(5-\sqrt{5})}\chi_{\mathbb{I}}(\tilde{q}^{1/n})\,.\end{split}$ (79)
Therefore, the symmetry resolved Rényi entropy for the $\rho$-th irreducible block is
$\begin{split}S_{n}(\mathcal{R}_{\rho})&=\frac{1}{1-n}\log\!\left(\frac{Z(q^{n},\mathcal{R}_{\rho})}{Z^{n}(q,\mathcal{R}_{\rho})}\right),\quad\text{with}\quad\frac{Z(q^{n},\mathcal{R}_{\rho})}{Z^{n}(q,\mathcal{R}_{\rho})}\xrightarrow{\tilde{q}\to 0}p_{\rho}^{\,n-1}\,\tilde{q}^{\big(n-\frac{1}{n}\big)\frac{c_{\rm eff}}{24}},\end{split}$ (80)
where the coefficients are $p_{\rho}=\tfrac{1}{3},\forall\rho$. Then it yields the following for SREE (22),
$\begin{split}{S}(\mathcal{R}_{\rho})&=\frac{c}{3}\log\!\left(\frac{L}{\epsilon}\right)+\log(g^{2})+\log\!\left(\frac{d_{\rho}}{\sum d_{\rho}}\right),\,\textrm{with,}\,\,g^{2}=3\sqrt{\frac{1}{30}(5-\sqrt{5})},\,d_{\rho}=1\\ &\to\frac{c}{3}\log\!\left(\frac{L}{\epsilon}\right)+\log(g^{2})+\log\!\left(\frac{1}{|G|}\right),\,|G|=\textrm{cardinality of $\mathbb{Z}_{3}$.}\end{split}$ (81)
The quantum dimensions for the three lines are $d_{\rho}=1$. In Capizzi:2021kys, the authors have obtained the same result using the form factor bootstrap method.
5 SREE in Non-Unitary Rational CFT
As we already discussed in the introduction, symmetry resolution has mainly been developed in unitary settings, i.e, for unitary fusion categories, modular data, and RCFTs. The non-unitary case remains largely unexplored. By bulk–boundary correspondence, unitary TQFTs admit unitary boundary RCFT Witten:1988hf; Moore:1988qv; Moore:1989vd; Dijkgraaf:1989pz; Verlinde:1988sn; Fuchs:2009iz, and extending this logic beyond unitarity is a natural question Gang:2023rei. Recent work shows that topological twists of certain $3d$ $\mathcal{N}=4$ rank-zero (S-fold) SCFTs yield non-unitary TQFTs with modular data generalizing the Haagerup–Izumi constructions Gang:2023ggt; Gang:2022kpe. Furthermore, modular data of non-unitary rational CFTs have been constructed using bulk-boundary correspondence and used the method of Bantay:2005vk; Bantay:2007zz [^13] For other interesting applications of this method please refer to Govindarajan:2025jlq; Govindarajan:2025rgh. to compute the characters [^14] One other method to compute characters of a rational CFT is known as MMS approach, by analyzing modular invariant linear differential equation (MLDE) Mathur:1988na; Mathur:1988gt; Naculich:1988xv. For more details, readers are further referred to Hampapura:2015cea; Das:2021uvd; Chandra:2018pjq. We begin our analysis by examining the simplest non-unitary RCFT, known as the Lee-Yang minimal model.
5.1 Lee-Yang Minimal Model $\mathcal{M}(5,2)$: a simplest example
The Lee-Yang minimal model is the simplest non-unitary rational CFT with central charge $c=-\frac{22}{5}$. It has two primaries with non-zero conformal dimension $h=-\frac{1}{5}$. This model possesses the Fibonacci fusion category symmetry, an example of non-invertible symmetry. The non-trivial fusion is given by
$\Phi_{2}\times\Phi_{2}=\Phi_{1}+\Phi_{2}$ (82)
where $\Phi_{1}$ corresponds to the vacuum state and $\Phi_{2}$ corresponds to a non-vacuum state, in this case, the non-vacuum state is the ground state $|\Omega\rangle$. For unitary theory, the vacuum state is the ground state, but not in non-unitary theory. It is a diagonal RCFT, and its modular S-matrix and characters are given by
[…]
$\chi_{1}(\tau)=q^{\frac{11}{60}}\left(1+q^{2}+q^{3}+\ldots\right)\,$, (84)
$\chi_{2}(\tau)=q^{-\frac{1}{60}}\left(1+q+q^{2}+\ldots\right)\,.$ (85)
We want to study this Fibonacci fusion category symmetry-resolved entanglement entropy for this non-unitary minimal model by leveraging the techniques of BCFT as discussed earlier. The weakly symmetric Cardy boundary state and the annulus partition function in the open-string channel are given by
$|\Phi_{2}\rangle=-i\left(1-\frac{2}{\sqrt{5}}\right)^{\frac{1}{4}}|1\rangle\rangle+\frac{\sqrt{5+3\sqrt{5}}}{\sqrt[4]{10\left(5+\sqrt{5}\right)}}|2\rangle\rangle,$ (86)
$Z(q)=\chi_{1}(q)+\chi_{2}(q)\,.$ (87)
The Fibonacci fusion category $\mathcal{C}=\{\Phi_{1},\Phi_{2}\}$ has two elements; there are two irreps. Therefore, the reduced density matrix has two diagonal blocks. The action of $\mathcal{C}$ on the primaries is given in the Table (4).
Defect line $\Phi_{1}$ $\Phi_{2}$
$\mathcal{L}_{\Phi_{1}}$ 11 11
$\mathcal{L}_{\Phi_{2}}$ $\frac{1-\sqrt{5}}{2}$ $\frac{1+\sqrt{5}}{2}$
Table 4: Action of the Fibonacci category on the primaries.
This action distinguishes two sectors. Therefore, the partition function is fully resolved, and in each block, there is a contribution from a single character. So, the symmetry-resolved partition functions are
$\begin{split}&Z(q^{n},\mathcal{R}_{1})=\frac{\chi_{1}(q^{n})}{(Z(q))^{n}},\,\,\,\,\,\,\,\,\,Z(q^{n},\mathcal{R}_{2})=\frac{\chi_{2}(q^{n})}{(Z(q))^{n}}\,.\end{split}$ (88)
Now, we can compute SREE. This is a non-unitary rational CFT; therefore, the leading contribution comes from the ground state rather than the vacuum. In this example, $|\Omega\rangle=|\Phi_{2}\rangle\rangle$ is a ground state. Therefore,
$\frac{Z(q^{n},\mathcal{R}_{1})}{Z^{n}(q,\mathcal{R}_{1})}\overset{\tilde{q}\to 0}{\sim}\left(\sqrt{\frac{5-\sqrt{5}}{10}}\right)^{1-n}\widetilde{q}^{\,(n-\frac{1}{n})\frac{c_{\rm eff}}{24}}$ (89)
and we get the SREE for the block $\mathcal{R}_{1}$
$S(\mathcal{R}_{1})=\frac{c_{\rm eff}}{3}\log\left(\frac{L}{\epsilon}\right)+\frac{1}{2}\log\frac{5-\sqrt{5}}{10}+\ldots$ (90)
where the constant term
$\frac{1}{2}\log\frac{5-\sqrt{5}}{10}=\log\left(\frac{d_{\mathcal{R}_{1}}}{\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}}\right)+\log(\langle\!\langle\Phi_{2}|\!|\Omega\rangle\!\langle\Omega|\!|\Phi_{2}\rangle\!\rangle).$ (91)
Here $d_{\mathcal{R}_{1}}=1$, $\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}=\frac{3+\sqrt{5}}{2}$, and $\log(\langle\!\langle\Phi_{2}|\!|\Omega\rangle\!\langle\Omega|\!|\Phi_{2}\rangle\!\rangle)=\log\left(\frac{5+3\sqrt{5}}{\sqrt{10(5+\sqrt{5})}}\right)$ corresponds to the Affleck-Ludwig boundary entropy Affleck:1991tk. Similarly, we have
$S(\mathcal{R}_{2})=\frac{c_{\rm eff}}{3}\log\left(\frac{L}{\epsilon}\right)+\frac{1}{2}\log\frac{5+\sqrt{5}}{10}+\ldots$ (92)
and the constant term
$\frac{1}{2}\log\frac{5+\sqrt{5}}{10}=\log\left(\frac{d_{\mathcal{R}_{2}}}{\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}}\right)+\log(\langle\!\langle\Phi_{2}|\!|\Omega\rangle\!\langle\Omega|\!|\Phi_{2}\rangle\!\rangle).$ (93)
where, $d_{\mathcal{R}_{2}}=\frac{1+\sqrt{5}}{2}$ is the quantum dimension of $\Phi_{2}$.
5.2 The Haagerup Rational CFT $\mathcal{R}_{\kappa=3}$
We now turn to one of the central topics of our paper, the symmetry resolution of generalized Haagerup RCFTs. We begin with the simplest example, $\mathcal{R}_{k=3}$, realized as the tensor product of the Virasoro minimal models $\mathcal{M}(2,5)$ and $\mathcal{M}(3,5)$ . The modular S-matrix is given by Gang:2023ggt,
$\begin{split}S=\left(\begin{array}[]{cccccccc}-a_{0}&a_{0}&a_{3}&-a_{3}&a_{2}&-a_{2}&a_{2}&-a_{2}\\ a_{0}&a_{0}&a_{3}&a_{3}&-a_{2}&-a_{2}&-a_{2}&-a_{2}\\ a_{3}&a_{3}&a_{0}&a_{0}&a_{2}&a_{2}&a_{2}&a_{2}\\ -a_{3}&a_{3}&a_{0}&-a_{0}&-a_{2}&a_{2}&-a_{2}&a_{2}\\ a_{2}&-a_{2}&a_{2}&-a_{2}&2a_{2}\cos\left(\frac{\pi}{5}\right)&2a_{2}\cos\left(\frac{2\pi}{5}\right)&2a_{2}\cos\left(\frac{3\pi}{5}\right)&2a_{2}\cos\left(\frac{4\pi}{5}\right)\\ -a_{2}&-a_{2}&a_{2}&a_{2}&2a_{2}\cos\left(\frac{2\pi}{5}\right)&2a_{2}\cos\left(\frac{4\pi}{5}\right)&2a_{2}\cos\left(\frac{6\pi}{5}\right)&2a_{2}\cos\left(\frac{8\pi}{5}\right)\\ a_{2}&-a_{2}&a_{2}&-a_{2}&2a_{2}\cos\left(\frac{3\pi}{5}\right)&2a_{2}\cos\left(\frac{6\pi}{5}\right)&2a_{2}\cos\left(\frac{9\pi}{5}\right)&2a_{2}\cos\left(\frac{12\pi}{5}\right)\\ -a_{2}&-a_{2}&a_{2}&a_{2}&2a_{2}\cos\left(\frac{4\pi}{5}\right)&2a_{2}\cos\left(\frac{8\pi}{5}\right)&2a_{2}\cos\left(\frac{12\pi}{5}\right)&2a_{2}\cos\left(\frac{16\pi}{5}\right)\\ \end{array}\right)\end{split}$ (94)
where, $a_{0}=\frac{1}{\sqrt{8(k+2)}}+\frac{1}{\sqrt{8(k-2)}},a_{1}=\frac{1}{\sqrt{2(k-2)}},a_{2}=\frac{1}{\sqrt{2(k+2)}},\,a_{3}=\frac{1}{\sqrt{8(k-2)}}-\frac{1}{\sqrt{8(k+2)}}$, with $k=3$. The full fusion can be immediately found by the Verlinde formula. However, by carefully inspecting the full fusion one can find that the following primaries $\Phi_{1},\Phi_{4},\,\Phi_{5},\,\Phi_{7}$ form a sub-fusion. The sub-fusion is given by,
$\begin{split}&\Phi_{4}\times\Phi_{4}=\Phi_{1}+\Phi_{4}+\Phi_{5}+\Phi_{7},\,\\ &\Phi_{5}\times\Phi_{4}=\Phi_{4}+\Phi_{7},\,\\ &\Phi_{5}\times\Phi_{5}=\Phi_{1}+\Phi_{5},\\ &\Phi_{7}\times\Phi_{4}=\Phi_{4}+\Phi_{5},\\ &\Phi_{7}\times\Phi_{5}=\Phi_{4},\\ &\Phi_{7}\times\Phi_{7}=\Phi_{1}+\Phi_{7}\,.\end{split}$ (95)
The sub-fusion category has four elements and, accordingly, four irreducible representations $\{\mathcal{R}_{i}\}_{i=1}^{6}$ . Consequently, the reduced density matrix of subsystem $A$, $\rho_{A}$, decomposes into four distinct (mutually orthogonal) blocks, each carrying a unique character. However, the attainable resolution depends on whether the defect lines act distinctly on the primaries: primaries with identical defect action remain in the same block. To proceed further, we investigate the action of (topological) defects on the primary fields. We carry out our analysis using the diagonal RCFT associated with the $\mathcal{R}_{k=3}$ modular data, where primaries and topological defects are in one-to-one correspondence. Equivalently, each defect $\mathcal{L}_{a}$ (labeled by a primary $a$) acts diagonally on primaries:
$\mathcal{L}_{a}\Phi_{i}\;=\;\lambda_{a}(i)\,\Phi_{i},\qquad\lambda_{a}(i)\;=\;\frac{S_{ai}}{S_{1i}},$ (96)
where $S$ is the modular $S$-matrix. Table (5) summarizes the action of defects on the primaries.
$\Phi_{1}$ $\Phi_{4}$ $\Phi_{5}$ $\Phi_{7}$
$\mathcal{L}_{1}$ 1 1 1 1
$\mathcal{L}_{4}$ $\frac{1}{2}\left(3-\sqrt{5}\right)$ $\frac{1}{2}\left(\sqrt{5}+3\right)$ 1-1 1-1
$\mathcal{L}_{5}$ $\frac{1}{2}\left(1-\sqrt{5}\right)$ $\frac{1}{2}\left(\sqrt{5}+1\right)$ $\frac{1}{2}\left(\sqrt{5}+1\right)$ $\frac{1}{2}\left(1-\sqrt{5}\right)$
$\mathcal{L}_{7}$ $\frac{1}{2}\left(1-\sqrt{5}\right)$ $\frac{1}{2}\left(\sqrt{5}+1\right)$ $\frac{1}{2}\left(1-\sqrt{5}\right)$ $\frac{1}{2}\left(\sqrt{5}+1\right)$
Table 5: Eigenvalues $\lambda_{\mathcal{L}_{a}}(i)$ for the action of defect $\mathcal{L}_{a}$ on primary $\Phi_{i}$ in a diagonal RCFT, with $\lambda_{a}(i)=\frac{S_{ai}}{S_{1i}}$.
In this non-unitary RCFT the state of lowest conformal weight (the “ground state” for thermodynamic purposes) is $|\Phi_{3}\rangle$, not the identity. Consequently, the quantum dimensions of the topological defect lines $\mathcal{L}_{a}$ are evaluated relative to $|\Phi_{3}\rangle$ via the modular $S$-matrix:
$d_{\mathcal{L}_{a}}\;=\;\frac{S_{a3}}{S_{13}}\,.$ (97)
For the defects of interest, we obtain
$d_{\mathcal{L}_{1}}$ $=1\,$, $d_{\mathcal{L}_{4}}$ $=\tfrac{1}{2}\!\left(\sqrt{5}+3\right)\,$, $d_{\mathcal{L}_{5}}$ $=\tfrac{1}{2}\!\left(\sqrt{5}+1\right)\,$, $d_{\mathcal{L}_{7}}$ $=\tfrac{1}{2}\!\left(\sqrt{5}+1\right)\,.$ (98)
To compute the annulus partition function, we choose the (weakly) symmetric boundary state $|\!|\,\Phi_{4}\rangle\!\rangle$. The closed–string channel amplitude is then
[…]
where $\tilde{q}$ is the modular parameter in the closed channel. The partition function in (99) can be computed by expanding the boundary Cardy states in terms of Ishibashi states as,
$|\!|{B}_{i}\rangle\!\rangle=\sum_{i}\frac{S_{ij}}{\sqrt{S_{1j}}}|j\rangle\rangle,\,\textrm{with},\,{B_{i}}=\Phi_{4},$ (100)
where $|j\rangle\rangle$’s are the Ishibashi states and using the following property of Ishibashi states: $\langle\langle i|\tilde{q}^{L_{0}-\frac{c}{24}}|j\rangle\rangle=\chi_{i}(\tilde{q})\delta_{ij}$. In the open string channel, the partition function takes the following form (16),
$\begin{split}Z(q)&=\chi_{1}(q)+\chi_{4}(q)+\chi_{5}(q)+\chi_{7}(q)\,.\end{split}$ (101)
As shown in the Table (5), the action of defects $\mathcal{L}_{a}$ is clearly different on the primaries, hence we have four irreducible blocks. The leading contribution to the partition function at $\tilde{q}\to 0$ limit is given by,
$\begin{split}&Z(q^{n},\Phi_{1})\sim\frac{5-\sqrt{5}}{10\sqrt{2}}\chi_{3}(\tilde{q}^{\frac{1}{n}}),\,\quad Z(q^{n},\Phi_{4})\sim\frac{1}{2}\sqrt{\frac{3}{5}+\frac{1}{\sqrt{5}}}\,\chi_{3}(\tilde{q}^{\frac{1}{n}}),\quad\,Z(q^{n},\Phi_{5})\sim\frac{1}{\sqrt{10}}\,\chi_{3}(\tilde{q}^{\frac{1}{n}}),\,\\ &\qquad\qquad\qquad\qquad\qquad\qquad Z(q^{n},\Phi_{7})\sim\frac{1}{\sqrt{10}}\chi_{3}(\tilde{q}^{\frac{1}{n}})\,.\end{split}$ (102)
Therefore, the SREE for the first irreducible block is given by,
$\begin{split}S(\mathcal{R}_{1})=\frac{c_{\textrm{eff}}}{3}\log\left(\frac{L}{\epsilon}\right)+\log\left(\frac{5-\sqrt{5}}{10\sqrt{2}}\right)+\cdots\end{split}$ (103)
Therefore the consistency condition from (20) is the following,
$\begin{split}\log\left(\frac{5-\sqrt{5}}{10\sqrt{2}}\right)\overset{?}{=}\log(g^{2})+\log\left(\frac{d_{\rho}\,N_{\rho B_{1}}^{B_{2}}}{d(B_{1}\otimes B_{2})}\right)\,.\end{split}$ (104)
Now computing the RHS independently, we find,
$g^{2}=|\langle\langle\Phi_{4}||\Omega\rangle\rangle|^{2}=\sqrt{\frac{2}{5}}+\frac{1}{\sqrt{2}},\,\,\log\left(\frac{d_{\rho}\,N_{\rho B_{1}}^{B_{2}}}{d(B_{1}\otimes B_{2})}\right)=\frac{1}{2}\left(7-3\sqrt{5}\right)\,.$ (105)
Putting (105) in (104), one can show that the equality in (20) (or more generally (23)) indeed holds. Note that we choose the boundary as weakly symmetric, for which we notice that $d(B_{1}\otimes B_{2})=\sum_{\rho}d_{\mathcal{L}_{\rho}}$, and the expression for the entanglement entropy reduces to (22) .
Next, we turn to the example of $\mathcal{R}_{k=4}$ Gang:2023ggt, where several subtleties arise. In particular, the relevant sub-fusion ring exhibits non-trivial multiplicities, and we find that the standard methods do not yield a complete symmetry resolution. We make a detailed comment on that.
5.3 The Haagerup Rational CFT $\mathcal{R}_{\kappa=4}$
Haagerup $\mathcal{R}_{\kappa=4}$ is also a non-unitary rational conformal field theory with ten primaries with central charge c=11c=-11 and conformal dimensions are $h_{i}=\{\frac{3}{2},\,-\frac{1}{2},\,0,\,-\frac{3}{8},\,-\frac{11}{24},\,-\frac{1}{3},\,-\frac{1}{8},\,\frac{1}{6},\,\frac{13}{24}\}$ Gang:2023ggt, which can also be realized as non-unitary super-Virasoro minimal model $S\mathcal{M}(2,12)$. The modular S-matrix by which theses characters transform under $\tau\to-\frac{1}{\tau}$ is given by Gang:2023ggt,
$S=\begin{pmatrix}a_{0}&a_{0}&a_{3}&a_{3}&-a_{1}&a_{2}&-a_{2}&a_{2}&-a_{2}&a_{2}\\ a_{0}&a_{0}&a_{3}&a_{3}&a_{1}&-a_{2}&-a_{2}&-a_{2}&-a_{2}&-a_{2}\\ a_{3}&a_{3}&a_{0}&a_{0}&a_{1}&a_{2}&a_{2}&a_{2}&a_{2}&a_{2}\\ a_{3}&a_{3}&a_{0}&a_{0}&-a_{1}&-a_{2}&a_{2}&-a_{2}&a_{2}&-a_{2}\\ -a_{1}&a_{1}&a_{1}&-a_{1}&2a_{1}\cos\left(\frac{\pi}{2}\right)&0&0&0&0&0\\ a_{2}&-a_{2}&a_{2}&-a_{2}&0&2a_{2}\cos\left(\frac{\pi}{6}\right)&2a_{2}\cos\left(\frac{2\pi}{6}\right)&2a_{2}\cos\left(\frac{3\pi}{6}\right)&2a_{2}\cos\left(\frac{4\pi}{6}\right)&2a_{2}\cos\left(\frac{5\pi}{6}\right)\\ -a_{2}&-a_{2}&a_{2}&a_{2}&0&2a_{2}\cos\left(\frac{2\pi}{6}\right)&2a_{2}\cos\left(\frac{4\pi}{6}\right)&2a_{2}\cos\left(\frac{6\pi}{6}\right)&2a_{2}\cos\left(\frac{8\pi}{6}\right)&2a_{2}\cos\left(\frac{10\pi}{6}\right)\\ a_{2}&-a_{2}&a_{2}&-a_{2}&0&2a_{2}\cos\left(\frac{3\pi}{6}\right)&2a_{2}\cos\left(\frac{6\pi}{6}\right)&2a_{2}\cos\left(\frac{9\pi}{6}\right)&2a_{2}\cos\left(\frac{12\pi}{6}\right)&2a_{2}\cos\left(\frac{15\pi}{6}\right)\\ -a_{2}&-a_{2}&a_{2}&a_{2}&0&2a_{2}\cos\left(\frac{4\pi}{6}\right)&2a_{2}\cos\left(\frac{8\pi}{6}\right)&2a_{2}\cos\left(\frac{12\pi}{6}\right)&2a_{2}\cos\left(\frac{16\pi}{6}\right)&2a_{2}\cos\left(\frac{20\pi}{6}\right)\\ a_{2}&-a_{2}&a_{2}&-a_{2}&0&2a_{2}\cos\left(\frac{5\pi}{6}\right)&2a_{2}\cos\left(\frac{10\pi}{6}\right)&2a_{2}\cos\left(\frac{15\pi}{6}\right)&2a_{2}\cos\left(\frac{20\pi}{6}\right)&2a_{2}\cos\left(\frac{25\pi}{6}\right)\end{pmatrix}$
where, $a_{0}=\frac{1}{\sqrt{8(\kappa-2)}}+\frac{1}{\sqrt{8(\kappa+2)}},\,a_{1}=\frac{1}{\sqrt{2(\kappa-2)}},\,a_{2}=\frac{1}{\sqrt{2(\kappa+2)}},\,a_{3}=\frac{1}{\sqrt{8(\kappa-2)}}-\frac{1}{\sqrt{8(\kappa+2)}},\,\kappa=4$. Therefore, we can compute the fusion rule of the primaries, and we have noticed that there is a sub-fusion algebra generated by $\mathcal{C}=\{\Phi_{1},\Phi_{2},\Phi_{3},\Phi_{4},\Phi_{7},\Phi_{9}\}$.
$\Phi_{2}\times\Phi_{2}$ $=\Phi_{1}\,$,
$\Phi_{3}\times\Phi_{2}$ $=\Phi_{4}\,$,
$\Phi_{3}\times\Phi_{3}$ $=\Phi_{1}+\Phi_{3}+\Phi_{4}+\Phi_{7}+\Phi_{9}\,$,
$\Phi_{4}\times\Phi_{2}$ $=\Phi_{3}\,$,
$\Phi_{4}\times\Phi_{3}$ $=\Phi_{2}+\Phi_{3}+\Phi_{4}+\Phi_{7}+\Phi_{9}\,$,
$\Phi_{4}\times\Phi_{4}$ $=\Phi_{1}+\Phi_{3}+\Phi_{4}+\Phi_{7}+\Phi_{9}\,$,
$\Phi_{5}\times\Phi_{2}$ $=\Phi_{5}$
[…]
We also notice that the boundary state $|\!|\,\Phi_{5}\rangle\!\rangle$ is weakly symmetric under this fusion category $\mathcal{C}$.
$\begin{split}&\mathcal{L}_{\Phi_{1}}|\!|\,\Phi_{5}\rangle\!\rangle=|\!|\,\Phi_{5}\rangle\!\rangle\,,\\ &\mathcal{L}_{\Phi_{2}}|\!|\,\Phi_{5}\rangle\!\rangle=|\!|\,\Phi_{5}\rangle\!\rangle\,,\\ &\mathcal{L}_{\Phi_{3}}|\!|\,\Phi_{5}\rangle\!\rangle=|\!|\,\Phi_{5}\rangle\!\rangle\oplus|\!|\,\Phi_{5}\rangle\!\rangle\oplus|\!|\,\Phi_{6}\rangle\!\rangle\oplus|\!|\,\Phi_{8}\rangle\!\rangle\oplus|\!|\,\Phi_{10}\rangle\!\rangle\,,\\ &\mathcal{L}_{\Phi_{4}}|\!|\,\Phi_{5}\rangle\!\rangle=|\!|\,\Phi_{5}\rangle\!\rangle\oplus|\!|\,\Phi_{5}\rangle\!\rangle\oplus|\!|\,\Phi_{6}\rangle\!\rangle\oplus|\!|\,\Phi_{8}\rangle\!\rangle\oplus|\!|\,\Phi_{10}\rangle\!\rangle\,,\\ &\mathcal{L}_{\Phi_{7}}|\!|\,\Phi_{5}\rangle\!\rangle=|\!|\,\Phi_{5}\rangle\!\rangle\oplus|\!|\,\Phi_{5}\rangle\!\rangle\oplus|\!|\,\Phi_{6}\rangle\!\rangle\oplus|\!|\,\Phi_{8}\rangle\!\rangle\oplus|\!|\,\Phi_{10}\rangle\!\rangle\,,\\ &\mathcal{L}_{\Phi_{9}}|\!|\,\Phi_{5}\rangle\!\rangle=|\!|\,\Phi_{5}\rangle\!\rangle\oplus|\!|\,\Phi_{5}\rangle\!\rangle\oplus|\!|\,\Phi_{6}\rangle\!\rangle\oplus|\!|\,\Phi_{8}\rangle\!\rangle\oplus|\!|\,\Phi_{10}\rangle\!\rangle\,.\end{split}$ (118)
The untwisted partition function in the closed string channel and open string channel are given respectively by,
$\begin{split}Z(q^{n})&=\frac{3}{3+\sqrt{3}}\,\chi_{1}(\tilde{q}^{\frac{1}{n}})+\frac{3}{3+\sqrt{3}}\,\chi_{2}(\tilde{q}^{\frac{1}{n}})+\frac{3}{3-\sqrt{3}}\,\chi_{3}(\tilde{q}^{\frac{1}{n}})+\frac{3}{3-\sqrt{3}}\,\chi_{4}(\tilde{q}^{\frac{1}{n}})\,,\\ &=\chi_{1}(q^{n})+\chi_{2}(q^{n})+2\chi_{3}(q^{n})+2\chi_{4}(q^{n})+\chi_{7}(q^{n})+\chi_{9}(q^{n})\,.\end{split}$ (119)
Note that the chosen weakly symmetric boundary state, $|\!|{B}\rangle\!\rangle=|\!|\,\Phi_{5}\rangle\!\rangle$ does not belong to the sub-fusion. However, from the fusion table, one can immediately see that ${B}\otimes{B}$ again generates all the elements of the sub-fusion. From Table (6), it is clear that the actions of the six defect lines are the same for the following blocks: $\{\Phi_{1},\,\Phi_{2}\},\,\{\Phi_{3},\,\Phi_{4}\},\,\{\Phi_{7},\Phi_{9}\}$. Therefore, there will be three blocks (reducible) to be resolved.
Defect line/Bulk primary $\Phi_{1}$ $\Phi_{2}$ $\Phi_{3}$ $\Phi_{4}$ $\Phi_{5}$ $\Phi_{6}$ $\Phi_{7}$ $\Phi_{8}$ $\Phi_{9}$ $\Phi_{10}$
$\mathcal{L}_{\Phi_{1}}$ 11 11 11 11 11 11 11 11 11 11
$\mathcal{L}_{\Phi_{2}}$ 11 11 11 11 1-1 1-1 11 1-1 11 1-1
$\mathcal{L}_{\Phi_{3}}$ $2-\sqrt{3}$ $2-\sqrt{3}$ $2+\sqrt{3}$ $2+\sqrt{3}$ 1-1 11 1-1 11 1-1 11
$\mathcal{L}_{\Phi_{4}}$ $2-\sqrt{3}$ $2-\sqrt{3}$ $2+\sqrt{3}$ $2+\sqrt{3}$ 11 1-1 1-1 1-1 1-1 1-1
$\mathcal{L}_{\Phi_{7}}$ $1-\sqrt{3}$ $1-\sqrt{3}$ $1+\sqrt{3}$ $1+\sqrt{3}$ 0 11 11 2-2 11 11
$\mathcal{L}_{\Phi_{9}}$ $1-\sqrt{3}$ $1-\sqrt{3}$ $1+\sqrt{3}$ $1+\sqrt{3}$ 0 1-1 11 22 11 1-1
Table 6: Action of the defect lines on primaries.
The leading contribution to the partition function at $\tilde{q}\to 0$ limit is given by,
[…]
The partially resolved entanglement entropy for each block is given by,
$\begin{split}&S(\mathcal{R}_{1})=\frac{c_{\textrm{eff}}}{3}\log\left(\frac{L}{\epsilon}\right)+\log\left(\frac{1}{6}\left(3-\sqrt{3}\right)\right)\,,\\ &S(\mathcal{R}_{2})=\frac{c_{\textrm{eff}}}{3}\log\left(\frac{L}{\epsilon}\right)+\log\left(1+\frac{1}{\sqrt{3}}\right)\,,\\ &S(\mathcal{R}_{3})=\frac{c_{\textrm{eff}}}{3}\log\left(\frac{L}{\epsilon}\right)+\log\left(\frac{1}{\sqrt{3}}\right)\,.\end{split}$ (121)
Now we ask the same question as before by making a comparison with (20),
$\begin{split}&\textrm{block-1}:\,\,\log\left(\frac{1}{6}\left(3-\sqrt{3}\right)\right)\overset{?}{=}\log(g^{2})+\log\left(\frac{d_{\rho}\,N_{\rho B_{1}}^{B_{2}}}{d(B_{1}\otimes B_{2})}\right),\,\rho\in\mathcal{R}_{1}\,,\\ &\textrm{block-2}:\,\,\log\left(1+\frac{1}{\sqrt{3}}\right)\overset{?}{=}\log(g^{2})+\log\left(\frac{d_{\rho}\,N_{\rho B_{1}}^{B_{2}}}{d(B_{1}\otimes B_{2})}\right),\,\rho\in\mathcal{R}_{2}\,,\\ &\textrm{block-3}:\log\left(\frac{1}{\sqrt{3}}\right)\overset{?}{=}\log(g^{2})+\log\left(\frac{d_{\rho}\,N_{\rho B_{1}}^{B_{2}}}{d(B_{1}\otimes B_{2})}\right),\,\rho\in\mathcal{R}_{3}\,.\end{split}$ (122)
Again, independent computation shows,
$\begin{split}&g^{2}=|\langle\!\langle\Phi_{5}|\!|\Omega\rangle|^{2}=\frac{1}{2}\left(\sqrt{3}+3\right),\,\\ &\frac{\sum_{\imath\in\rho}d_{\imath}\,N_{\imath B_{1}}^{B_{2}}}{d(B_{1}\otimes B_{2})}=\left\{\frac{2}{\left(\sqrt{3}+3\right)^{2}},\frac{2}{3},\,\left(\frac{1}{\sqrt{3}}-\frac{1}{3}\right)\right\},\quad\forall\imath\in\rho=1,2,3\quad(\textrm{respectively})\,.\end{split}$ (123)
Substituting (123) into (122) confirms that the equality of (20) remains valid. However, (22) does not apply here: the multiplicity-free assumption is violated, the fusion exhibits nontrivial multiplicities, and also the resulting blocks are not made up from single characters. Apart from that, one may notice that the defect lines discussed above do not fully resolve the symmetry sector. Can we further refine the decomposition into smaller irreducible blocks? The answer is to some extent-yes, thanks to Heymann:2024vvf.
Prescription. Refinement via an auxiliary defect. Introduce an additional defect line associated with the boundary state $|\!|\,\Phi_{5}\rangle\!\rangle$, denoted by $\mathcal{L}_{\Phi_{5}}$, and examine its action on the primaries of the chosen subcategory. If this defect acts differently on the primaries within a single block $\mathcal{R}_{i}$, that is, if the images/eigenvalues under $\mathcal{L}_{\Phi_{5}}$ are not identical, then $\mathcal{R}_{i}$ splits into two irreducible blocks; otherwise, no further refinement occurs. Applying this prescription to our case, the action of $\mathcal{L}_{\Phi_{5}}$ separates the primaries in the first two blocks, thereby splitting each into two irreducible blocks, while it fails to further decompose the third block. Consequently, we obtain five blocks in total—four irreducible and one remaining reducible. A straightforward calculation, in the same spirit as above, once again confirms (20).
5.4 The Haagerup Rational CFT $\mathcal{R}_{\kappa=5}$
Finally, we consider another example of a generalized Haagerup fusion category, namely $\mathcal{R}_{\kappa=5}$ Gang:2023ggt. The theory possesses six integrable highest-weight representations labeled by spins, $j=0,\tfrac{1}{2},1,\tfrac{3}{2},2,\tfrac{5}{2}.$ Below, we discuss the fusion of primaries in the fusion category. We again find that a sub-fusion exists and is given by,
$\Phi_{4}\times\Phi_{4}$ $=\Phi_{1}+\Phi_{4}+\Phi_{5}+\Phi_{7}+\Phi_{9}+\Phi_{11}\,$,
$\Phi_{4}\times\Phi_{5}$ $=\Phi_{4}+\mathbf{2}\Phi_{5}+\Phi_{7}+\Phi_{9}+\Phi_{11}\,$,
$\Phi_{4}\times\Phi_{7}$ $=\Phi_{4}+\Phi_{5}+\Phi_{9}+\Phi_{11}\,$,
$\Phi_{4}\times\Phi_{9}$ $=\Phi_{4}+\Phi_{5}+\Phi_{7}+\Phi_{11}\,$,
$\Phi_{4}\times\Phi_{11}$ $=\Phi_{4}+\Phi_{5}+\Phi_{7}+\Phi_{9}\,$,
$\Phi_{5}\times\Phi_{5}$ $=\Phi_{1}+\mathbf{2}\Phi_{4}+\mathbf{2}\Phi_{5}+\Phi_{7}+\Phi_{9}+\Phi_{11}\,$,
$\Phi_{5}\times\Phi_{7}$ $=\Phi_{4}+\Phi_{5}+\Phi_{7}+\Phi_{9}+\Phi_{11}\,$,
$\Phi_{5}\times\Phi_{9}$ $=\Phi_{4}+\Phi_{5}+\Phi_{7}+\Phi_{9}+\Phi_{11}\,$,
$\Phi_{5}\times\Phi_{11}$ $=\Phi_{4}+\Phi_{5}+\Phi_{7}+\Phi_{9}+\Phi_{11}\,$,
$\Phi_{7}\times\Phi_{7}$ $=\Phi_{1}+\Phi_{5}+\Phi_{7}+\Phi_{9}\,$,
$\Phi_{7}\times\Phi_{9}$ $=\Phi_{4}+\Phi_{5}+\Phi_{7}\,$,
$\Phi_{7}\times\Phi_{11}$ $=\Phi_{4}+\Phi_{5}+\Phi_{11}\,$,
$\Phi_{9}\times\Phi_{9}$ $=\Phi_{1}+\Phi_{5}+\Phi_{9}+\Phi_{11}\,$,
$\Phi_{9}\times\Phi_{11}$ $=\Phi_{4}+\Phi_{5}+\Phi_{9}\,$,
$\Phi_{11}\times\Phi_{11}$ $=\Phi_{1}+\Phi_{5}+\Phi_{7}+\Phi_{11}\,.$
Defect line/Bulk primary $\Phi_{1}$ $\Phi_{4}$ $\Phi_{5}$ $\Phi_{7}$ $\Phi_{9}$ $\Phi_{11}$
$\mathcal{L}_{\Phi_{1}}$ 11 11 11 11 11 11
$\mathcal{L}_{\Phi_{4}}$ $\frac{5-\sqrt{21}}{2}$ $\frac{5+\sqrt{21}}{2}$ 11 1-1 1-1 1-1
$\mathcal{L}_{\Phi_{5}}$ $\frac{7-\sqrt{21}}{2}$ $\frac{7+\sqrt{21}}{2}$ 1-1 0 0 0
$\mathcal{L}_{\Phi_{7}}$ $\frac{3-\sqrt{21}}{2}$ $\frac{3+\sqrt{21}}{2}$ 0 $2\cos(\frac{\pi}{7})$ $2\sin(\frac{\pi}{14})$ $2\sin(\frac{3\pi}{14})$
$\mathcal{L}_{\Phi_{9}}$ $\frac{3-\sqrt{21}}{2}$ $\frac{3+\sqrt{21}}{2}$ 0 $2\sin(\frac{\pi}{14})$ $2\sin(\frac{3\pi}{14})$ $2\cos(\frac{\pi}{7})$
$\mathcal{L}_{\Phi_{11}}$ $\frac{3-\sqrt{21}}{2}$ $\frac{3+\sqrt{21}}{2}$ 0 $2\sin(\frac{3\pi}{14})$ $2\cos(\frac{\pi}{7})$ $2\sin(\frac{\pi}{14})$
Table 7: Action of the defect lines on primaries.
The sub-fusion category contains six elements, $\mathcal{C}=\{\Phi_{1},\Phi_{4},\Phi_{5},\Phi_{7},\Phi_{9},\Phi_{11}\}$, so it has six irreducible representations, $\mathcal{R}_{i}$. We can see from the Table (7) that the action of $\mathcal{C}$ on all the primaries is distinct. Therefore, we conclude that a single character will contribute in each representation/block in the reduced density matrix $\rho_{A}$, namely, if we take $||\Phi_{4}\rangle\rangle$ as a weakly symmetric boundary state, we have the following partition function and symmetry-resolved partition functions in the open-string channel respectively,
$Z(q)=\chi_{1}(q)+\chi_{4}(q)+\chi_{5}(q)+\chi_{7}(q)+\chi_{9}(q)+\chi_{11}(q)$ (124)
and,
$\begin{split}&Z(q^{n},\mathcal{R}_{1})=\frac{\chi_{1}(q^{n})}{(Z(q))^{n}}\,,\,\,\,\,\,\,Z(q^{n},\mathcal{R}_{2})=\frac{\chi_{4}(q^{n})}{(Z(q))^{n}}\,,\,\,\,\,\,\,\,Z(q^{n},\mathcal{R}_{3})=\frac{\chi_{5}(q^{n})}{(Z(q))^{n}}\,,\\ &Z(q^{n},\mathcal{R}_{4})=\frac{\chi_{7}(q^{n})}{(Z(q))^{n}}\,,\,\,\,\,\,\,\,Z(q^{n},\mathcal{R}_{5})=\frac{\chi_{9}(q^{n})}{(Z(q))^{n}}\,,\,\,\,\,\,\,\,Z(q^{n},\mathcal{R}_{6})=\frac{\chi_{11}(q^{n})}{(Z(q))^{n}}\,.\end{split}$ (125)
This is a non-unitary rational CFT; therefore leading contribution comes from the ground state rather than the vacuum in $overset{\tilde{q}\to 0}\,.$ In this example, $|\Phi_{3}\rangle$ is a ground state. Finally, we get the following expressions for the SREE for the different blocks.
$S_{n}(\mathcal{R}_{1})=\frac{c_{\rm eff}}{3}\log\left(\frac{L}{\epsilon}\right)+\log\frac{1}{2}\left(\frac{1}{\sqrt{6}}-\frac{1}{\sqrt{14}}\right)+\ldots\,.$ (126)
where the constant term
$\log\frac{1}{2}\left(\frac{1}{\sqrt{6}}-\frac{1}{\sqrt{14}}\right)=\log\left(\frac{d_{\mathcal{R}_{1}}}{\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}}\right)+\log(\langle\!\langle\Phi_{4}|\!|\Omega\rangle\langle\Omega|\!|\Phi_{4}\rangle\!\rangle)\,.$ (127)
Here $d_{\mathcal{R}_{1}}=1$, $\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}=\frac{23+5\sqrt{21}}{2}$, and $\log(\langle\Phi_{4}|\Omega\rangle\langle\Omega|\Phi_{4}\rangle)=\frac{1}{2}\log\left(\frac{55}{42}+2\sqrt{\frac{3}{7}}\right)$ corresponds to the Affleck-Ludwig boundary entropy Affleck:1991tk. Similarly, we have
$S_{n}(\mathcal{R}_{2})=\frac{c_{\rm eff}}{6}\left(\frac{n+1}{n}\right)\log\left(\frac{L}{\epsilon}\right)+\log\frac{1}{2}\left(\sqrt{\frac{5+\sqrt{21}}{21}}\right)+\ldots\,.$ (128)
and the constant term
$\log\frac{1}{2}\left(\sqrt{\frac{5+\sqrt{21}}{21}}\right)=\log\left(\frac{d_{\mathcal{R}_{2}}}{\sum\limits_{\mathcal{R}_{i}}d_{\mathcal{R}_{i}}}\right)+\log(\langle\!\langle\Phi_{4}|\!|\Omega\rangle\langle\Omega|\!|\Phi_{4}\rangle\!\rangle)$ (129)
where, $d_{\mathcal{R}_{2}}=\frac{5+\sqrt{21}}{2}$ is the quantum dimension of $\mathcal{L}_{\Phi_{4}}\,.$
6 Summary and Conclusion
In this work, we have investigated the symmetry-resolved entanglement entropy (SREE) in $(1+1)$-dimensional critical systems possessing categorical symmetries. Our interest has been in a broad class of rational conformal field theories (RCFTs), including examples that are (1) non-unitary, (2) non-diagonal, and (3) exhibit multiplicities in their fusion rules, for which symmetry resolution is substantially more intricate than in the conventional group-like or unitary diagonal cases. The general motivation for our study comes from the ongoing effort to understand how entanglement structures encode internal and categorical symmetry data in quantum field theories. In particular, recent developments in Symmetry Topological Field Theory (SymTFT) Bhardwaj:2024qiv; Bhardwaj:2025piv; Bhardwaj:2023ayw; Schafer-Nameki:2025fiy; GarciaEtxebarria:2024jfv; Bhardwaj:2023bbf; Apruzzi:2025hvs; Bhardwaj:2024igy and the study of defect networks provide a powerful framework to interpret charges and sectors in RCFTs. However, explicit computations of SREE in non-unitary and non-diagonal theories, especially those with non-group-like symmetry categories, have remained limited. This work aims to fill that gap.
Summary of Results.
Our analysis and main results may be summarized as follows:
  • General Framework for Categorical Symmetry Resolution. We first discussed the role of boundary conditions in symmetry resolution and introduced a general prescription, inspired by the SymTFT construction of Choi:2024wfm, for computing symmetry-resolved entanglement entropies in arbitrary categorical charge sectors. This formalism applies not only to weakly and strongly symmetric boundary conditions, but also to the more subtle cloaking boundary conditions, and it naturally accommodates fusion categories with multiplicities. We demonstrated that the resulting formula is universally applicable across a wide range of RCFTs, irrespective of unitarity or diagonality.
  • Replica Computations, Universal Scaling, and Boundary entropy. We carried out explicit replica computations of the von Neumann entropy associated with each symmetry sector, identifying both the universal leading contributions and the subleading corrections associated with boundary entropy. For several unitary and non-unitary models, we evaluated the SREE via first principles and made a connection with the generalized Affleck–Ludwig boundary entropy formula. We began with the Tambara–Yamagami (TY) fusion category structure in the Ising CFT, where the absence of weakly/strongly symmetric boundary conditions necessitates cloaking boundary states. We also analyze the SREE for the tetracritical Ising model. Interestingly, in this case, we found that, there were two natural choices of sub-fusion categories. For the larger sub-fusion category we find that no weakly symmetric boundary state exists, but we can still perform the resolution using an appropriate cloaking boundary state. In contrast, for the smaller sub-fusion category, a weakly symmetric boundary state was available. We subsequently turned to non-diagonal RCFTs. In the $SU(2)_{10}$ theory with $E_{6}$ modular invariant, we showed that symmetry resolution with respect to the $\tfrac{1}{2}E_{6}$ fusion ring fails: although one may choose a weakly symmetric boundary state $|\!|Y\rangle\!\rangle$, the open-channel spectrum contains primaries that do not appear in the $E_{6}$ fusion ring. Therefore, the resolution is ambiguous and no consistent sector decomposition exists. By contrast, for the three-state Potts model, we demonstrated that symmetry resolution with respect to the $\mathbb{Z}_{3}$ symmetry is possible, provided one selects an appropriate weakly symmetric boundary condition.
  • Non-Unitary Symmetry Resolution and Generalized Haagerup RCFTs. We then studied the symmetry resolution in non-unitary settings. As a starting point, we computed the SREE for the Lee–Yang minimal model. Following this, we addressed generalized Haagerup RCFTs, whose modular data encode non-group-like categorical symmetries and which may not necessarily admit a well-defined VOA realization. This makes the implementation of boundary-based SREE calculations particularly subtle. We reviewed the modular and categorical data of the Haagerup family $\mathcal{R}_{k}$, discussed the structure of simple objects and fusion channels, and emphasized the non-invertible symmetry interpretation. By carefully identifying admissible Cardy boundary states, we explicitly computed the symmetry-resolved entanglement entropies for $\mathcal{R}_{k=3,4,5}$. Our findings show that with an appropriate choice of weakly symmetric boundary conditions, full symmetry-sector resolution is achievable for $\mathcal{R}_{3}$ and $\mathcal{R}_{5}$. However, for $\mathcal{R}_{4}$, a complete resolution is not possible when restricting to defect subfusion sectors alone. We demonstrated that introducing boundary defects allows one to improve and, in some cases, complete the resolution.
Outlook and open problem There are several interesting future directions possible. We list a few of them below. It is highly compelling to consider extending these computations beyond the realm of rational conformal field theories (RCFTs). This naturally raises several fundamental questions, for instance, whether the existence of a vertex operator algebra (VOA) is a necessary condition for the existence of a corresponding symmetry topological field theory (SymTFT)? A particularly promising direction is to refine and generalize the existing framework so as to address the challenges of symmetry resolution in non-diagonal RCFTs, where modular invariants and fusion structures become considerably more intricate.
In addition to this, our current analysis has primarily focused on line operators, which correspond to 0-form defects. A natural next step would be to explore the broader landscape of higher-form defects, such as condensation or domain-wall defects, which play a crucial role in the categorical and geometric understanding of symmetry breaking and topological phase transitions. Another significant avenue of investigation lies in the explicit computation of the anomaly inflow action. Determining this action would allow one to identify and characterize the electric and magnetic defect lines, along with their associated bulk linkings. Such an analysis could shed light on the interplay between boundary symmetries and bulk topological data, ultimately contributing to a deeper understanding of the correspondence between conformal and topological field theories.
Acknowledgments
We would like to thank Brandon Rayhaun for detailed comments on the previous version of the manuscript. A.B and J.S would also like to thank Arpit Das and Thomas Quella for discussions on related topics and comments on some parts of the draft, as well as making us aware of the upcoming work Das:2025xyz. J.S is indebted to Suresh Govindarajan, Chethan N. Gowdigere, and Akhila S. for useful discussions. S.G (PMRF ID: 1702711) and S.P (PMRF ID: 1703278) is supported by the Prime Minister’s Research Fellowship of the Government of India. S.G would like to thank Institut des Hautes Études Scientifiques (IHES), France, for their kind hospitality during the course of the work. A.B thanks the participants of the BIRS-CMI workshop “Quantum Gravity and Information Theory: Modern Developments” (https://www.birs.ca/events/2025/5-day-workshops/25w5386) held at CMI, Chennai, during 9-14th November 2025 where part of the work was presented by J.S. J.S would like to thank the organizers of the BIRS-CMI workshop 2025 for giving warm hospitality. J.S is supported by the project CRG/2023/001120 and OTH/SERB/13451 (50:50 Postdoctoral scheme of IIT Gandhinagar). A.B is supported by the Core Research Grant (CRG/2023/001120) from the Anusandhan National Research Foundation, India. A.B also acknowledges the associateship program of the Indian Academy of Science, Bengaluru.
Appendix A Another non-unitary minimal model $\mathcal{M}(7,2)$
Very recently, there has been growing interest in the minimal model $\mathcal{M}(7,2)$ in the context of RG flows and non-invertible symmetries Ambrosino:2025yug; Katsevich:2025ojk; Blakeney:2025ext; Fukusumi:2025xrj. $\mathcal{M}(7,2)$ is a diagonal non-unitary Virasoro minimal model with central charge $c=-\frac{68}{7}$ and non-vanishing conformal weights are $\{-\frac{2}{7},-\frac{3}{7}\}$. It has three Virasoro primaries with non-trivial fusion
$\Phi_{2}\times\Phi_{2}$ $=$ $\Phi_{1}+\Phi_{3}\,$, (130)
$\Phi_{2}\times\Phi_{3}$ $=$ $\Phi_{2}+\Phi_{3}\,$, (131)
$\Phi_{3}\times\Phi_{3}$ $=$ $\Phi_{1}+\Phi_{2}+\Phi_{3}\,.$ (132)
[…]
Defect line $\Phi_{1}$ $\Phi_{2}$ $\Phi_{3}$
$\mathcal{L}_{\Phi_{1}}$ 11 11 11
[…]
Table 8: Action of the $\textrm{psu}(2)_{5}$ category on the primaries of $\mathcal{M}(7,2)$.
It turns out that the Cardy boundary state $|\!|\Phi_{3}\rangle\!\rangle$ is a weakly symmetric under the above fusion category, known as $\textrm{psu}(2)_{5}$ (It is a self-dual, multiplicity-free fusion category). This Cardy boundary state in terms of Ishibashi states is given by
[…]
The annulus partition function in the closed and open-string channels with boundary state $|\!|\Phi_{3}\rangle\!\rangle$ is given by
[…]
$=$ $\chi_{1}(q)+\chi_{2}(q)+\chi_{3}(q)\,.$ (135)
From the action of the defect lines on the primaries listed in Table (8), we can write the resolved-partition function in the three sectors, namely,
$Z(\mathcal{R}_{1},q^{n})$ $=$ $\frac{\chi_{1}(q^{n})}{(Z(q))^{n}}\,$, (136)
$Z(\mathcal{R}_{2},q^{n})$ $=$ $\frac{\chi_{2}(q^{n})}{(Z(q))^{n}}\,$, (137)
$Z(\mathcal{R}_{3},q^{n})$ $=$ $\frac{\chi_{3}(q^{n})}{(Z(q))^{n}}\,.$ (138)
$\mathcal{M}(7,2)$ is a non-unitary rational CFT; therefore, the leading contribution comes from the ground state rather than the vacuum. In this example, $|\Omega\rangle=|\Phi_{3}\rangle$ is a ground state. Therefore, the symmetry-resolved $n$th Rényi entropy is
[…]
where the constant term
[…]
and the constant terms are given by
[…]